### v2

Abstract
Scientific machine learning and the advent of the Physics-Informed Neural Network (PINN) have shown high potential in their ability to solve complex differential equations. One example is the use of PINNs to solve the gravity field modeling problem — learning convenient representations of the gravitational potential from position and acceleration data. These PINN gravity models, or PINN-GMs, have demonstrated advantages in model compactness, robustness to noise, and sample efficiency when compared to popular alternatives; however, further investigation has revealed various failure modes for these and other machine learning gravity models which this manuscript aims to address. Specifically, this paper introduces the third generation Physics-Informed Neural Network Gravity Model (PINN-GM-III) which includes design changes that solve the problems of feature divergence, bias towards low-altitude samples, numerical instability, and extrapolation error. Six evaluation metrics are proposed to expose these past pitfalls and illustrate the PINN-GM-III’s robustness to them. This study concludes by evaluating the PINN-GM-III modeling accuracy on a heterogeneous density asteroid, and comparing its performance to other analytic and machine learning gravity models.
keywords:
Scientific Machine Learning, Physics Informed Neural Networks, Astrodynamics, Gravity Modeling
1 Introduction
Nearly all problems in astrodynamics involve the force of gravity. Be it in trajectory optimization, spacecraft rendezvous, orbit determination, or other problems, gravity often plays a significant — if not dominant — role in the system dynamics. The mere ubiquity of this force is a testament to its significance, yet despite this, there exist surprisingly few ways to represent this force to high accuracy. The construction of high-fidelity gravity models is henceforth referred to as the gravity modeling problem, and currently there exists no universally adopted solution. Some gravity models perform well for large planetary bodies, but break down when modeling objects with exotic geometries like asteroids and comets. Other models can handle irregular shapes but require assumptions and come with high computational cost. The one commonality is that all existing gravity models come with their own unique pitfalls that prevent standardization across the community. Consequently, researchers are continuing to design new solutions to the gravity modeling problem with hopes of one day finding a universal model. One encouraging vein of research explores the use of neural networks and scientific machine learning to learn gravity models free of these past difficulties.
Figure 1: PINN-GMs offer high-accuracy, low-cost gravity solutions to be deployed across a variety of applications.
Scientific machine learning uses neural networks to generate high-fidelity models of, and solutions to, complex differential equations cuomoScientificMachineLearning2022 ; karniadakisPhysicsinformedMachineLearning2021 . Physics-informed neural networks (PINNs) are one class of model available to solve such problems. PINNs are neural networks trained in a manner that intrinsically respects relevant differential and physics-based constraints. This compliance, or “physics-informing”, is achieved by augmenting the network cost function with said constraints, such that the learned model is penalized for violating any of the underlying physics. Through this design change, PINNs have been shown to achieve higher accuracy with less data than their traditional neural network counterparts raissiPhysicsinformedNeuralNetworks2019 .
Recently, researchers have proposed the use of PINNs as well as other machine learning models to solve the gravity modeling problem martinEarthAndMoon2022 ; martinSmallBodies2022 . These models have been shown to produce high-accuracy solutions in both large- and small-body settings under specific training conditions. While these models demonstrate early promise, closer inspection reveals that there remain regimes in which these machine learning models perform unreliably. This paper exposes these shortcomings and proposes solutions to them through multiple design changes to the underlying machine learning architecture. Specifically, this paper introduces the third generation of the PINN gravity model, or PINN-GM, which includes design changes that solve the problems of feature divergence, extrapolation error, numerical instability, bias towards low-altitude samples, and incompliant boundary conditions.
2 Background
Gravity field models can be categorized into two groups: analytic and numerical. Analytic models are derived from first principles and yield closed-form equations for the gravitational potential, whereas numerical models are constructed in a data-driven manner that come without an explicit equation. Each model has its own set of advantages and drawbacks, and the choice of which model to use is often dictated by application. The following section briefly surveys the available gravity models and their corresponding pros and cons.
2.1 Analytic Models
Spherical Harmonics Model
In the 1900s, spherical harmonic basis functions were proposed to represent high-order perturbations in the Earth’s gravity field brillouinEquationsAuxDerivees1933 . These harmonics can be superimposed to produce the spherical harmonic gravity model through:
$U(r)=\frac{\mu}{r}\sum_{l=0}^{l}\sum_{m=0}^{l}\bigg{(}\frac{R}{r}\bigg{)}^{l}P_{l,m}\left(\sin\phi\right)\big{[}C_{l,m}\cos(m\lambda)+S_{l,m}\sin(m\lambda)\big{]}$ (1)
where $r$ is the distance to the test point, $\mu$ is the gravitational parameter of the body, $R$ is the circumscribing radius, $l$ and $m$ are the degree and order of the model, $C_{l,m}$ and $S_{l,m}$ are the Stokes coefficients, $\lambda$ and $\phi$ are the longitude and latitude, and $P_{l,m}$ are the associated Legendre polynomials kaulaTheorySatelliteGeodesy1966 . The spherical harmonic gravity model is commonly used to represent the fields of large planetary bodies like the Earth pavlisPreliminaryGravitationalModel2005 , the Moon goossensGlobalDegreeOrder2016 , and Mars genovaSeasonalStaticGravity2016a , as they are among the most efficient models for capturing planetary oblateness — the largest gravitational perturbation found on these large, rotating bodies. Using only a single coefficient, $C_{2,0}$, spherical harmonics can succinctly capture this important gravitational feature and its effects on spacecraft motion.
While these models are effective at representing planetary oblateness, they struggle to model the remaining gravitational perturbations like mountain ranges, tectonic plate boundaries, and craters. These discontinuous features are notoriously hard to represent using periodic basis functions — requiring the superposition of hundreds of thousands of high-frequency harmonics to overcome the 3D equivalent of Gibbs phenomenon hewittGibbsWilbrahamPhenomenonEpisode1979 . These higher frequencies introduce an $\mathcal{O}(n^{2})$ computational and memory cost martinGPGPUImplementationPines2020 , and the regression of these harmonics is especially difficult due to their rapidly fading observability. High-order spherical harmonic models also can diverge when evaluated within the sphere that bounds all mass elements (Brillouin sphere) due to the $(R/r)^{l}$ term in the series expansion. While these effects are negligible for near-spherical planets or moons, they can become problematic in small-body settings where objects can exhibit highly non-spherical geometries wernerExteriorGravitationPolyhedron1997 .
Polyhedral Gravity Model
The polyhedral gravity model is a popular alternative in these small-body settings, offering a solution that maintains validity down to the surface of any object regardless of shape. These models require a preexisting shape model of the asteroid or comet — a collection of triangular facets and vertices which captures its geometry — from which an analytic acceleration can be computed through:
$\nabla U=-G\sigma\sum_{e\in\text{edges}}\mathbf{E}_{e}\cdot\mathbf{r}_{e}\cdot L_{e}+G\sigma\sum_{f\in\text{facets}}\mathbf{F}_{f}\cdot\mathbf{r}_{f}\cdot\omega_{f}$ (2)
where $G$ is the gravitational constant, $\sigma$ is the density of the body, $\mathbf{E}_{e}$ is an edge dyad, $\mathbf{r}_{e}$ is the position vector between the center of the edge and the test point, $L_{e}$ is an analog to the potential contribution of the edge, $\mathbf{F}_{f}$ is the face normal dyad, $\mathbf{r}_{f}$ is the distance between the face normal and the test point, and $\omega_{f}$ is an analog to the potential contribution of the face wernerExteriorGravitationPolyhedron1997 .
While the polyhedral gravity model avoids divergence within the Brillouin sphere, it comes with its own challenges. First, this gravity model is expensive when evaluated on high-resolution shape models — requiring intensive summation loops over all vertices and facets to compute the acceleration at a single test point. The model also assumes that the body’s density is known a priori. While a constant density assumption is often used, literature has shown that such assumption is not necessarily valid scheeresEstimatingAsteroidDensity2000 ; scheeresHeterogeneousMassDistribution2020 .
Mascon Gravity Model
In contrast to the analytically derived spherical harmonic and polyhedral models, the mascon gravity models instead approximates the gravitational potential using a collection of point mass elements, known as mascons mullerMasconsLunarMass1968 . These masses are distributed within the body and summed to form a global approximation of the gravity field through Equation 3:
$\bm{a}(\bm{r})=\sum_{k=0}^{N}\mu_{k}\frac{\bm{r}-\bm{r}_{k}}{\lVert\bm{r}-\bm{r}_{k}\rVert^{3}}$ (3)
where $\mu_{k}$ and $\bm{r}_{k}$ are the estimated gravitational parameter and position for the $k$-th mascon.
While the mascon model offers an efficient alternative to the polyhedral model, it becomes inaccurate when evaluated near individual mascons tardivelLimitsMasconsApproximation2016 . Hybrid mascon models offer a more accurate alternative — representing each mascon with a low fidelity spherical harmonic model — but this incorporates additional complexity in their regression and remains prone to the same challenges of the traditional mascon and spherical harmonic models wittickMixedmodelGravityRepresentations2019 .
Other Analytic Models
The ellipsoidal harmonic model follows a similar approach to that of spherical harmonics but uses ellipsoidal harmonic basis functions instead romainEllipsoidalHarmonicExpansions2001 . This yields a smaller region in which the model could diverge; however, it still struggles to model discontinuity with its periodic basis functions. The interior spherical harmonic model inverts the spherical harmonic formulation and can model a local region whose boundary intersects only one point on the surface of the body takahashiSurfaceGravityFields2013 . This model maintains stability down to that single point making it valuable for precise landing operations; however, the solution becomes invalid on any other point on the surface and outside of the corresponding local sphere. Finally, the interior spherical Bessel gravity model expands upon the interior model but uses Bessel functions rather than spherical harmonics to achieve a wider region of validity. This model comes with added analytic complexity and can also struggle to capture discontinuous features efficiently takahashiSmallBodySurface2014 .
2.2 Machine Learning Models
As an alternative to analytic gravity models, recent efforts explore the use of machine learning to regress models of complex gravity fields in a data-driven manner. Unlike analytic approaches, machine learning models are generally free of assumptions about the body they are modeling and have no analytic limitations — e.g. divergence in the bounding sphere, required shape models, etc. These models have historically required large volumes of training data; however once trained, they can offer high-accuracy predictions at comparatively low computational cost. Table 1 summarizes key accuracy and training metrics for recently reported machine learning gravity models, and a specific discussion for each model is provided below.
Model Parameters Training Data Avg. Error [%] Valid Globally
GP gaoEfficientGravityField2019 12,960,000 3,600 1.5%
NNs chengRealtimeControlFueloptimal2019 1,300,000 800,000 0.35%
ELMs furfaroModelingIrregularSmall2020 350,000 768,000 1-10%
GeodesyNet izzoGeodesyIrregularSmall2022 80,800 500,000 0.36%
PINN-GM-III 2,211 4,096 0.30%
Table 1: Machine Learning Gravity Model Statistics – See Appendix C
Gaussian Processes
Gaussian processes, or GPs, are non-parametric models that are fit by specifying a prior distribution over functions, and updating that prior based on observed data. This requires the user to first specify some kernel function which measures similarity between data and then compute a covariance matrix between all data pairs using that function. Once computed, the covariance matrix is inverted and used to evaluate the mean and uncertainty of the learned function at a test point.
GPs were proposed to solve the gravity modeling problem in 2019 gaoEfficientGravityField2019 , regressing a mapping between position and acceleration training data. GPs are advantageous because they provide a probabilistic estimate of the uncertainty in the model’s prediction; however these models do not scale well to large data sets. The GP’s covariance matrix is built from the training data and scales as $\mathcal{O}(n^{2})$, where $n$ is the size of the training data set, and the computational complexity of the matrix inversion scales as $\mathcal{O}(n^{3})$. This scaling makes it impractical to fit GPs using large quantities of data, intrinsically limiting its utility and performance.
Extreme Learning Machines
Extreme learning machines (ELMs) have also been proposed to predict gravitational accelerations furfaroModelingIrregularSmall2020 . ELMs are single layer neural networks fit by randomly initializing the weights from the inputs to the hidden layer, and then computing the optimal weights to the output layer using least-squares regression. Explicitly the ELM minimizes the mean-square error loss function:
$L(\bm{\theta})=\frac{1}{N}\sum_{i=0}^{N}\lVert\hat{\bm{y}}_{i}(\bm{x}_{i}|\bm{\theta})-\bm{y}_{i}\rVert^{2}$ (4)
where $\hat{y}_{i}(\bm{x_{i}}|\bm{\theta})$ is the machine learning model prediction at input $\bm{x}_{i}$ with trainable model parameters $\bm{\theta}$ huangExtremeLearningMachine2006 .
ELMs are advantageous because they can model non-linear functions, and they only require a single training iteration. However, depending on the width of the ELM and the amount of training data used, these models can be prone to memory and computational difficulties due to the large $\mathcal{O}(n^{3})$ matrix inversion used in the least squares solution. Iterative chunking strategies have been proposed to remedy this issue, though this makes the relative advantage between ELMs and traditional neural networks less apparent. Moreover, these models are data intensive, such that past solutions have required hundreds of thousands of training data to produce high-accuracy models.
Neural Networks
Neural networks are similar to ELMs, but they are typically multi-layer and fit using gradient decent and backpropagation rather than least squares. These models iteratively update their parameters using small batches of training data combined with various optimization algorithms kingmaAdamMethodStochastic2014 ; dozatIncorporatingNesterovMomentum ; nocedalNumericalOptimization2006 . Neural networks have also been proposed as a candidate gravity model chengRealtimeOptimalControl2020 . When juxtaposed with ELMs, neural networks’ primary benefit is their lack of required matrix inverse; however, this comes at the cost of iterative training which can take long periods of time. Neural networks share similar drawbacks to ELMs, requiring large quantities of data and are prone to extrapolation error when evaluated outside of the bounds of the training data.
Physics-Informed Neural Networks
Physics-informed neural networks, or PINNs, are yet another model proposed to solve the gravity modeling problem martinEarthAndMoon2022 ; martinSmallBodies2022 . These physics-informed models increase sample efficiency over traditional networks by incorporating differential constraints into the loss function of the neural network. These constraints limit the set of learnable functions to only those that comply with the underlying physics raissiPhysicsinformedNeuralNetworks2019 . The first PINN gravity model, or PINN-GM, uses the known differential equation $\nabla U=\bm{a}$ to define the loss function:
[…]
where $\hat{U}$ is the potential learned by the network, which can then be differentiated via automatic differentiation baydinAutomaticDifferentiationMachine2018 to compute the corresponding acceleration. The inclusion of these physics constraints, combined with network design modifications, have shown that PINN-GMs can yield solutions that maintain similar accuracy to their predecessors while using orders of magnitude fewer parameters and data. These models have also demonstrated enhanced robustness to uncertainty in the training data as a result of their physics-informed constraints martinSmallBodies2022 .
GeodesyNet
Finally, in 2022 Reference 25 introduced GeodesyNets as a candidate solution to the gravity modeling problem. GeodesyNets are neural density fields — close relatives to the popular neural radiance fields mildenhallNeRFRepresentingScenes2022 — which uses neural networks to learn a density function for every point in a 3D volume. Once learned, these density fields can be integrated to yield a gravitational acceleration or potential. This work is an exciting new class of machine learning model capable of achieving high accuracy; albeit the cost of the numerical integration of densities make it cumbersome for simulation or on-board purposes. These models also have more computationally expensive training processes than other neural network approaches.
2.3 Current Challenges
As these machine learning gravity models become more mainstream, they warrant further scrutiny. While these models can yield accurate solutions under ideal training and test cases, further investigation reveal that these models are not yet universally robust. Rather, there exist common problems that exist for the majority of these models that have yet to be identified or addressed. Specifically, these models are prone to extrapolation error and numerical instabilities, and they are also relatively brittle to sparse and noisy data conditions. If the community strives to develop a universally valid machine learning gravity model, these challenges need to be addressed.
In this manuscript, we aim to expose these various failure cases and introduce design modifications to the underlying machine learning architectures that can increase the their robustness and generalizability. To accomplish this, we introduce the third generation PINN gravity model, or PINN-GM-III, which includes a variety of design changes that improve modeling accuracy and robustness across a wide set of training and test cases. These modifications, and the failure modes they eliminate, are discussed at length in Section 3. While these modifications are applied specifically to the PINN-GM, it should be noted that some can be applied to other machine learning gravity models as well. Details of the specific PINN modifications and their impact on performance are supplied below.
3 PINN-GM-III
The training process for the original PINN-GM can be found in Ref. martinSmallBodies2022 and is briefly summarized here for convenience. To begin, the neural network requires a set of $N$ position and acceleration vectors sampled around a celestial body:
$\bar{\bm{x}}=\begin{bmatrix}\bm{x}_{1}\\ \bm{x}_{2}\\ \vdots\\ \bm{x}_{N}\end{bmatrix}\quad\bar{\bm{a}}=\begin{bmatrix}\bm{a}_{1}\\ \bm{a}_{2}\\ \vdots\\ \bm{a}_{N}\end{bmatrix}$ (6)
This data can be collected in one of two ways: 1) It can be generated synthetically from pre-existing high-fidelity gravity models like EGM-2008 for Earth pavlisDevelopmentEvaluationEarth2012 or a high-fidelity polyhedral model for asteroids gaskellGaskellErosShape2021 , or 2) the data can be estimated on-board without any prior knowledge of the field. The latter can be accomplished using estimation techniques like dynamic model compensation leonardGravityErrorCompensation2013 to simultaneously estimate spacecraft position and accelerations in-situ as is shown in Ref. martinPINNFilter2022 .
Figure 2: PINN-GM Generation III with new modifications highlighted in green
Once the data is gathered, a neural network is fit to learn a mapping between these data. The neural network takes the form:
$\hat{y}_{\bm{\theta}}(\bm{x})=W^{(L)}\cdot\sigma(W^{(L-1)}\cdot\ldots\cdot\sigma(W^{(1)}\cdot x+b^{(1)})+b^{(L-1)})+b^{(L)}$ (7)
where $\hat{y}_{\bm{\theta}}$ is the output of the neural network parameterized by weights and biases ${{W},{b}}\in\bm{\theta}$. These parameters are found in the 0th to L-th hidden layers of the network. Successive non-linear transformations of the network input $\bm{x}$ are applied via the activation function, $\sigma$, to yield $\hat{y}_{\bm{\theta}}(\bm{x})$.
For PINN-GMs, the output of the neural network corresponds to the predicted gravitational potential, $\hat{U}$, at position $\bm{x}$. That potential is then differentiated to produce an acceleration vector which is used in the network’s physics-informed loss function:
[…]
The network is iteratively trained using stochastic gradient descent which estimates the gradient of the loss with respect the network parameters, $\nabla_{\bm{\theta}}L$. For every iteration, the parameters are updated via:
$\bm{\theta}^{\prime}\rightarrow\bm{\theta}-\eta\nabla_{\bm{\theta}}L$ (9)
where $\eta$ is a user-defined learning rate.
The second generation PINN-GM, or PINN-GM-II, enhances this training process by augmenting the loss function with additional constraints to increase robustness to noise. In addition, better input features are selected to increase sample efficiency. Despite these improvements, closer inspection has revealed that the PINN-GM-II has failure cases that remain unaddressed. While these pitfalls are explicitly exposed within the PINN-GM-II, they are general to other published machine learning gravity models to the best of the authors’ knowledge. The subsequent sections outline these pitfalls and propose architectural changes to the underlying model architecture to address them. Collectively, these changes form the third generation PINN gravity model, or PINN-GM-III. The modifications are illustrated in Figure 2 and discussed in detail in Sections 3.1 through 3.5, with their combined effect demonstrated in Section 3.6.
Algorithm 1 PINN-GM-III algorithm
1:Collect training data ($\bm{x},\bm{a}$) from:
(a) a pre-existing model
(b) online state estimates martinPINNFilter2022
2:Non-dimensionalize the training data $\triangleright$ App. A
3:Convert to 5D non-singular spherical coordinates ($r_{i}$, $r_{e}$, $s,t,u$) $\triangleright$ Sec. 3.1
4:Propagate through the neural network
5:Output proxy potential $U_{\text{NN}}$triangleright Sec. 3.3
6:Scale proxy potential into true potential $\hat{U}_{\text{NN}}$triangleright Sec. 3.3
7:(Optional) Fuse with weighted low-fidelity potential $\hat{U}_{\text{LF}}$triangleright Sec. 3.5
8:Enforce boundary conditions on the network potential $\hat{U}$triangleright Sec. 3.4
9:Autodifferentiate (AD) potential to produce acceleration $\hat{\bm{a}}$
10:if Training then
11:    Sum acceleration percent error and absolute error to form loss $\triangleright$ Sec. 3.2
12:    Compute gradients of loss function
13:    Update network parameters
14:end if
3.1 Feature Engineering
The first pitfall of past machine learning gravity models is the choice of inefficient and divergent input features. Cartesian position coordinates are most regularly used as inputs for these models; however, these coordinates are prone to two major drawbacks. First, they span the set of all real numbers, and the networks require data sampled across this entire domain to learn robust solutions for any possible value of $\bm{x}$. The PINN-GM-II first identified this problem and proposed a conversion to 4D non-singular spherical coordinates $(r,s,t,u)$ where $s,t,$ and $u$ are the sine of the angle between the test point and the cartesian axes. While the domain for $s,t,$ and $u$ was reduced to $[-1,1]$ which increases the model’s sample efficiency, the radial coordinate, $r$, remains prone to the second major drawback: feature divergence.
When evaluating these numerical models at points far from the body, the radial coordinate remains unbounded and can introduce numerical instabilities that cause the model to diverge. While the radial coordinate can be converted to $1/r$, this instead introduces instability for test points near the surface of the body. In both cases, the radial feature can have magnitudes greater than one which will cause many activation functions to prematurely saturate and decrease the learning efficiency of the model goodfellowDeepLearning2016 .
This pitfall can be addressed through a simple design modification where the Cartesian position coordinates are converted into a 5D spherical coordinate description of $(r_{i},r_{e},s,t,u)$ where $r_{e}$ and $r_{i}$ are two proxies of the test point radius, $r$, defined as
$r_{i}=\begin{cases}r&\quad r\in[0,R]\\ 1&\quad r\in[R,\infty)\end{cases}\quad\text{and}\quad r_{e}=\begin{cases}1&\quad r\in[0,R]\\ \frac{1}{r}&\quad r\in[R,\infty)\end{cases}$ (10)
Using this convention, the network will maintain the desired sample efficiency of past approaches while also ensuring that its input features can never diverge regardless of the location of any training or test point. This modification constitutes the first design change of the PINN-GM-III.
3.2 Modified Loss Function to Account for High-Altitude Samples
The second pitfall of past machine learning gravity models arises from their default loss functions which inadvertently decrease modeling accuracy at high-altitudes. Specifically, most machine learning gravity models use an absolute error loss function — e.g. mean squared error (MSE), root mean squared error (RMS). For the original PINNs, this was captured through the norm of the differenced acceleration vectors:
[…]
While these loss functions will successfully minimize the most flagrant residuals between the predicted and true acceleration vectors; they induce an undesirable effect at high-altitudes. Because accelerations produced near the surface of a celestial body have much larger magnitudes than the accelerations produced at high-altitudes, low-altitude errors will always appear disproportionately large compared to any errors at high-altitude. As a consequence, loss functions that minimize absolute errors will always prioritize low-altitudes samples, even if the high-altitude predictions are more erroneous in a relative sense.
Fortunately, this design flaw can be trivially remedied by augmenting the loss function with an term that captures relative error, like mean percent error:
[…]
The joining of these relative and absolute error loss terms eliminates this altitude sensitivity and constitutes the second design change of the PINN-GM-III.
3.3 Improve Numerics by Learning a Proxy to the Potential
Another common problem of past machine learning gravity models is numerical clipping. These models use the same matrix operations (i.e. the neural network) to predict both very large and very small values of the potential depending on the test point’s altitude. Unfortunately, these scales can lead to difficult numerics. When the altitude exceeds some critical threshold, the predictions grow too small, and the network prematurely clips their value to zero. This is problematic for both model training and inference — prematurely capping the maximum altitude for which these models are usable.
The PINN-GM-III addresses this problem by instead learning a more numerically favorable proxy to the potential, $U_{\text{NN}}$, defined as:
$U_{\text{NN}}=U*n(r);\quad n(r)=\begin{cases}1\quad r<R\\ r\quad r>R\\ \end{cases}$ (13)
where $U$ is the true potential and $n(r)$ is an altitude-dependent scaling function. This scaling function leverages the fact that potentials decay according to an inverse power-law outside the Brillouin sphere. By learning a potential normalized to $1/r$, the neural network predictions will not decay to unrecoverably small values but will instead remain bounded and centered about a non-dimensionalized value of $\mu$. This scaling eliminates numerical clipping to first order, granting stable numerics out to infinity. Importantly, once the network produces its numerically stable value for $U_{\text{NN}}$, the model will then explicitly transform that value into the true potential $\hat{U}_{\text{NN}}$ by dividing the output by the known scaling function through:
$\hat{U}_{\text{NN}}=\frac{U_{\text{NN}}}{n(r)}$ (14)
3.4 Enforcing Boundary Conditions via Transition Function to Avoid Extrapolation Error
The largest identified weakness of past machine learning gravity models is their extrapolation error. When tested beyond the bounds of their training data, machine learning models remain unconstrained and often diverge. The PINN-GM-III solves this problem by introducing a transition function that forces the model to seamlessly blend its network solution into a known analytic boundary condition outside the bounds of the training data through:
$\hat{U}(r)$ $=w_{\text{NN}}(r)\hat{U}_{\text{NN}}(r)+w_{\text{BC}}(r)U_{\text{BC}}(r)$ (15)
where $\hat{U}_{\text{NN}}$ is the predicted gravitational potential, $U_{\text{BC}}$ is the potential at the known boundary condition, and $w_{\text{NN}}$ and $w_{\text{BC}}$ are the altitude dependent weights for each defined as:
$w_{\text{BC}}(r,k,r_{\text{ref}})$ $=H(r,k,r_{\text{ref}})$ (16)
$w_{\text{NN}}(r,k,r_{\text{ref}})$ $=1-H(r,k,r_{\text{ref}})$ (17)
$H(r)$ is a smoothing function defined as
$H(r,k,r_{\text{ref}})=\frac{1+\tanh(k(r-r_{\text{ref}}))}{2}$ (18)
where $r$ is the distance to the test point, $r_{\text{ref}}$ is a reference radius, and $k$ is a smoothing parameter which controls the sharpness of the transition.
Equation (15) takes inspiration from Reference 38 which introduced the idea that PINNs can enforce physics compliance through other mechanisms than terms in their loss function. Explicitly, Reference 38 proposed the use of smooth, differentiable forms of the Heaviside function placed at the boundary to forcibly transition neural network outputs towards the known values.
In the case of gravity modeling, the most obvious boundary condition exists in the limit as $r\rightarrow\infty$, where the potential decays to zero. Setting $U_{\text{BC}}=0$ and $r_{\text{ref}}=\infty$ in Equation (15), however, is not practical, as it demands that the neural network must learn a model of the potential for the entire domain $r\in[0,\infty)$. A more useful choice is to leverage insights from the spherical harmonic gravity model and recognize that high frequency components of the gravitational potential decay to zero more quickly than the point mass contribution at high altitudes — i.e.
[…]
as $r\rightarrow\infty$. Therefore, $U_{\text{BC}}(r)$ can be set to $\frac{\mu}{r}$ assuming $r\gg R$. Therefore the PINN-GM-III sets $U_{\text{BC}}=\frac{\mu}{r}+f(r)$ in Equation (15), where $f(r)$ are any higher order terms in the spherical harmonic gravity model that the user knows a priori and wishes to leave as part of the boundary condition.
For completeness, the recommended value for $r_{\text{ref}}$ is the maximum altitude of the training data, and the recommended value for the transition coefficient is $k=0.5$. These choices ensure that the model will quickly transition to the boundary condition outside the training data, but without applying the transition too rapidly such that it changes the gradient of the potential and accidentally induces acceleration errors.
3.5 Leveraging Preexisting Gravity Information into PINN-GM Solution
The final design change of the PINN-GM-III proposes a mechanism for incorporating past analytic models into the network solution. Explicitly, analytic models remain popular for good reason, as they are often very good at representing certain parts of the gravity field — e.g. a point mass model captures first-order dynamics with a single parameter and and spherical harmonics can represent planetary oblateness with only $C_{2,0}$. Rather than abandoning these models and requiring the network to relearn these prominent behaviors and features, it would be far more convenient to incorporate these models into the learned solutions.
To accomplish this, the PINN-GM-III proposes modifying Equation (15) to include an analytic model term in the prediction:
$\hat{U}(r)=w_{\text{NN}}\big{(}U_{\text{LF}}(r)+U_{\text{NN}}(r)\big{)}+w_{\text{BC}}U_{\text{LF}}(r)$ (20)
where $U_{\text{LF}}$ refers to the known, low-fidelity analytic model such as $U_{\text{LF}}(r)=\frac{\mu}{r}+U_{J_{2}}(r)$. By incorporating these low-order models, the network can exclusively focus its modeling efforts on capturing high-order perturbations from these models.
Figure 3: Visualization of the various weighting factors applied to the neural network potential, boundary condition potential, and the low-fidelity potential.
(a) Error inside the training bounds
(b) Error outside the training bounds
Figure 4: Acceleration percent error as a function of altitude after sequentially applying each of the proposed PINN III modifications.
It should be noted that the inclusion of analytic models should be performed carefully. In the case of small-body gravity modeling, where the geometries may vastly differ from a point mass or low-fidelity spherical harmonic approximation, the incorporation of past analytic models can induce unnecessary error at low altitudes. To ensure that these analytic models are only used where appropriate, the hyperbolic tangent fusing function, $H(x,r,k)$, is reused to dynamically weight the analytic solution through:
$\hat{U}_{\text{LF}}(r)=w_{\text{LF}}U_{\text{LF}}(r)$ (21)
where $\hat{U}_{\text{LF}}$ is the weighted low fidelity analytic model with $w_{\text{LF}}=H(r,R^{*},k^{*})$, $R^{*}=1+e$, and $k=0.5$, where $e$ is the eccentricity of the body computed via $e=\sqrt{1-b^{2}/a^{2}}$. This ensures that at high altitude, where the analytic approximation is most accurate, it fully contributes to the final learned solution, but at lower altitudes, the analytic contribution has a reduced weight to the final solution. A visualization of the various weighting functions used in the PINN-GM-III is shown in Figure 3, highlighting where the different parts of the model are fully activated and deactivated.
3.6 Visualizing Modifications’ Effects on Model Performance
To visualize how the proposed modifications affect model accuracy, six PINN-GM are trained. The first model corresponds with the original PINN-GM-II, and the subsequent models sequentially add one of the modification discussed above. The PINN-GM are trained with a 5,000 point dataset uniformly distributed between the surface of the asteroid Eros to an altitude of 15R. The corresponding acceleration error of each model is reported within the training distribution in Figure 4(a) and outside the training distribution in Figure 4(b).
Figure 4(a) illustrates that these design modifications consistently improve the modeling accuracy of the PINN within the bounds of the training distribution. The PINN-GM-II had errors approaching 1% at the high 15R altitudes; however this decreases to 0.2%, 0.05%, 0.01%, 0.03%, and 0.001% for each of the respective modifications. Notably the transition to the new loss function (II) does decrease performance near the surface, but this is to be expected given the redistribution of modeling priorities that now balance accuracy at both low- and high-altitudes.
Figure 4(b) illustrates the stabilizing effect of these modifications on extrapolation error — quantifying prediction error outside of the training bounds out to 100R. As is shown in blue, the PINN-GM-II entirely diverges after leaving the bounds of the training data at 15R; however, the added design modifications stabilize performance and prevent divergence. As was expected, the error in the high-altitude limit hits a numerical plateau when only the features (I) and percent loss (II) modifications are included; however, when the proxy potential (III) is used, that plateau lowers by over an order-of-magnitude. The boundary condition modification (IV) eliminates this plateau — albeit inducing a slight penalty near the transition point due to changes in the gradient — however, the addition of the low-fidelity point mass potential (V) eliminates this penalty and further reduces the error within the bounds of the training data. It should be noted that there exists a fundamental numerical limit to the percent error at approximately 0.001% where the differences between model prediction and the ground truth exists at the limit of machine precision.
4 Benchmarking Suite
While the proposed modifications yield improvements as presented in Section 3.6, it remains important to characterize these gravity models from many different perspectives. The lack of comprehensive and standard performance benchmarks for these models are part of the reason these failure cases had not yet been identified. To eliminate this possibility moving forward and to further scrutinize the proposed PINN-GM-III, six new evaluation metrics are proposed. These metrics characterize gravity model accuracy within and beyond the training distribution, aiming to provide both coarse and fine measures of performance in different orbital regimes. These metrics are explained in detail in Section 4.1 and are used to characterize the performance of a trained PINN-GM-III on a heterogeneous density asteroid in Section 4.2.
4.1 Metrics
Planes Metric
The first accuracy metric assesses the mean percent error of the predicted acceleration vector along the three cartesian planes (XY, XZ, YZ) extended between [-5R, 5R] where R is the radius of the body. The field is evaluated on a 200x200 grid of points along each plane, and the average percent error is computed as
$P=\frac{1}{N}\sum_{i=1}^{N}\frac{\lVert\bm{a}_{\text{true}}-\bm{a}_{\text{PINN}}\rVert}{\lVert\bm{a}_{\text{true}}\rVert}\times 100$ (22)
This metric is intended to provide a coarse measure of model performance across a wide range of operational regimes.
Generalization Metrics
The second, third, and fourth metrics investigate the generalization of the model across a range of altitudes both within and beyond the training bounds. Explicitly, the mean acceleration error is evaluated as a function of altitude, and divided into three testing regimes: interior, exterior, and extrapolation. The interior metric assesses error within the bounding sphere of radius R. The exterior metric investigates the error out to the maximum altitude of the training dataset. Finally, the extrapolation metric measures the error exclusively outside the training dataset — specifically reaching altitudes 10 times larger than the maximum altitude represented in the training set. For every unit of radius, 500 samples are distributed uniformly in altitude to produce the test set.
Surface Metric
The fifth metric evaluates the mean acceleration error across all facets on a shape model of a celestial body, if available. This metric is used to characterize model performance at the most complex region of the field.
Trajectory Metric
The sixth metric evaluates the time-averaged position error of a trajectory propagated by the regressed model and the true trajectory of a spacecraft in a 24-hour low-altitude polar orbit about a rotating celestial body. Time-averaged error is used as it ensures that the error is monotonically increasing. To compute this value, the instantaneous position error $\Delta x(t)$ must be computed through
$\Delta x(t)=\lVert\underbrace{\bm{x}(t)}_{\text{True Pos.}}-\underbrace{\bm{\hat{x}}(t)}_{\text{Propagated Pos.}}\rVert$ (23)
from which the time-averaged error, $\mathcal{S}$, can be computed using numerical integration via:
$\mathcal{S}=\frac{1}{T}\int_{0}^{T}\Delta x(t)\text{d}t$ (24)
4.2 Experiment
These metrics are used to evaluate of a PINN-GM-III trained on data generated from a synthetic, heterogeneous density asteroid modeled after 433-Eros. Heterogeneous asteroids provide an especially challenging scenario for gravity models, as their internal density distributions are not directly observable. Some asteroids contain over- and under-dense regions within their interior, or may have been formed by two asteroids merging together. While some heuristic methods have been proposed to estimate these asteroids’ internal densities takahashiMorphologyDrivenDensity2014 which can be used in heterogeneous forms of the polyhedral model takahashiMorphologyDrivenDensity2014 ; the more common practice is to simply proceed with a constant density assumption.
To induce this heterogeneous density body, two small mass inhomogeneities are placed inside the asteroid 433-Eros. In one hemisphere, a mass element is added, and in the other hemisphere, a mass element is removed. Each mass element contains 10% of the total mass of the asteroid, and they are symmetrically displaced along the x-axis by 0.5R (see Figure 5(a)). The gravitational contributions of these mass elements are superimposed onto the gravity field of a constant density polyhedral model to form the simulated ground truth. This choice emulates the gravity field of a single body formed by two merged asteroids of different characteristic densities. The choice to make each mass element ±$\pm 10$% of the total mass is motivated based on literature with similar candidate density distributions kanamaruEstimationInteriorDensity2019 .
Using this heterogeneous density model, 90,000 position and acceleration data are sampled uniformly around the body from 0-10R. An additional 200,700 points are sampled on the surface of the asteroid — corresponding to a data point on every facet of the asteroid shape model. Together these approximately $300,000$ data points constitute a “best case” training set for the PINN-GM-III. These conditions are purposefully used to evaluate the upper-bound performance of the model. The next section will explore model performance under more realistic and stressful data conditions.
Using this best case data, a PINN-GM-II and a PINN-GM-III are trained. Both networks have six hidden layers with 32 nodes per layer which corresponds to approximately 5,300 learnable parameters. The default hyperparameters used to train these models can be found in Appendix A alongside the ablation studies used to select them in Appendix B. Once trained, these PINN-GMs are evaluated using the aforementioned metrics and compared to a constant-density polyhedral model as shown in Figure 6.
(a) Heterogeneous density asteroid
(b) Constant density polyhedral error
Figure 5: Heterogeneous density distributions found within asteroids can cause standard models and assumptions to break down.
PINN-GM-III Performance
Generalization
Planes
Surface
Trajectory
Figure 6: All proposed metrics evaluated for the constant density polyhedral model, the PINN-GM-II, and the PINN-GM-III.
The generalization metrics are shown at the top of Figure 6, highlighting model performance across the three altitude regimes. The constant density polyhedral model produces the highest error both on the interior ($<$R) and exterior (1-10R) metrics with errors averaging around 10% within Brillouin sphere and above 1% error outside of it. In the extrapolation regime, the polyhedral model error slowly decreases as the model tends towards a point mass approximation. In comparison, the PINN-GM-II produced average errors of 0.5% inside the bounding sphere, and 1% outside the bounding sphere, with errors slowly increasing at higher altitudes as a result of its original absolute error loss function. In the extrapolation regime, the PINN-GM-II entirely diverges for the reasons discussed in Section 3. In contrast to these models, the PINN-GM-III maintains lower errors across the interior and exterior distributions, averaging less than 0.5% error within the bounding sphere and approximately 0.005% error outside of it. As the model transitions into the extrapolation regime, the Heaviside function kicks in to enforce the point mass boundary condition — helping the network converge to the same limit as the polyhedral model.
The planes and surface metrics are shown in the middle rows of Figure 6. The polyhedral gravity model produced 6.5% average acceleration error on the planes, the PINN-GM-II reduces this to 0.4%, and the PINN-GM-III to 0.07%. The errors are largest near the surface for models, averaging at 23% for the constant density polyhedral model, 0.2% for the PINN-GM-II, and 0.18% for the PINN-GM-III.
Finally, the trajectory experiment is shown in the bottom row of Figure 6. For this experiment, the orbit is defined by the initial conditions $\{a,e,i,\omega,\Omega,M\}=\{32\text{ km},0.1,90^{\circ},0^{\circ},0^{\circ},0^{\circ}\}$, and the asteroid is rotating at $\omega_{0}=0.00073$ degrees per second along the z-axis. The figure on the left shows the instantaneous position error over the 24 hour integration period which, when time-averaged over the entire trajectory, yields average errors of 2,270 meters of error for the polyhedral model, 363 m for the PINN-GM-II, and 38 m for the PINN-GM-III.
Under these ideal data conditions, the PINN-GM-III achieves better performance on each benchmark when compared to the constant density polyhedral model and the prior PINN-GM-II. While these results are encouraging, the robustness of these models still needs to be tested under more realistic and stressful data conditions. Moreover it remains important that the PINN-GM-III is compared to other popular gravity models beyond just previous generation PINN-GMs. To accomplish this, the next section explores what happens to these models and other models under less ideal data conditions.
5 Comparative Study
In this section, a comparative study is performed to evaluate the PINN-GM-III against other popular gravity models. Explicitly, a point mass (PM), spherical harmonic (SH), mascon, polyhedral, extreme learning machine (ELM), traditional neural network (TNN), PINN-GM-I, and PINN-GM-II model are each regressed and then evaluated using the aforementioned metrics. Each model is fit eight times, permuting the data quantity, data uncertainty, and the size of each model. [^1] Note: The point mass model only has four permutations, as it does not have a “large” option and the polyhedral model only has two permutations, large and small. This is because polyhedral models are regressed in an entirely unrelated fashion to the other models — relying primarily on image data — and their sensitivity to data quantity and quality can not be adequately compared through this proposed experiment. Therefore, only the effect of model size is considered, and the performance can be considered an upper-bound. These conditions purposefully test model robustness under more stressful data conditions and different model sizes.
The dataset used for the comparison is generated by sampling position and acceleration data between 0-10R about the heterogeneous density asteroid discussed in Section 4. For the data sparse case, $N=500$ pairs are sampled. For the data rich case, $N=50,000$ pairs are sampled. These datasets have two configurations: noisy or noiseless. The noiseless configuration assumes ideal conditions in which the acceleration vectors are perfectly observable and have no error, whereas the noisy configuration perturbs every acceleration vector in a random direction by 10% of the acceleration magnitude via
$\bm{\tilde{a}}_{j}=\bm{a}_{j}+0.1\lVert\bm{a}_{j}\rVert\hat{\bm{u}_{j}}$
where $\hat{\bm{u}_{j}}$ is randomly sampled from the unit sphere. This error is chosen as a purposefully exaggerated test case to determine which of these models remain reliable under more stressful mission conditions.
In addition to training each model on these four datasets, this study also explores how the model size impacts performance. To do this, each model is tested at a small (S) and large (L) parametric capacity. These labels correspond to the total number of parameters used by each model—e.g. Stokes coefficients for a spherical harmonic model, facets and vertices in a shape model, or total weights and biases in a neural network. Each small model contain approximately 250 parameters, whereas the large models have approximately 30,000 parameters. Together, these varying data and parametric conditions provide insight into which models are capable of maintaining competitive performance in data sparse and low-memory regimes. Importantly, each model has slightly different regression procedures and model size calculations which are discussed in Appendix D.
N Error (%) Score Planes (%) Interior (%) Exterior (%) Extrap. (%) Surface (%) Traj. (km)
Model
PINN III S 50000 0 29 0.36 2.6 0.060 0.13 18 0.056
PINN III L 50000 0 31 1.2E-2 0.11 1.2E-3 1.1 3.3 8.4E-3
MASCONS L 50000 0 41 1.2E-2 0.13 5.4E-4 3.4E-4 33 2.7
PINN III S 500 0 64 1.5 8.6 0.35 0.27 31 1.1
PINN II S 50000 0 70 0.58 1.7 0.63 1.5E5 13 0.27
PINN II L 50000 0 73 1.4E-2 0.059 1.4E-2 1.4E15 2.8 9.6E-3
PINN III L 50000 10 76 3.4 3.7 2.0 0.28 13 6.3
MASCONS L 500 0 84 0.74 11 2.1E-4 9.8E-8 310 2.7
PINN III S 50000 10 92 3.9 8.3 2.0 0.34 35 3.0
MASCONS L 50000 10 94 1.3 9.3 0.16 0.15 630 2.4
POLY S 94 6.2 16 3.3 0.34 23 1.9
PINN II S 50000 10 95 2.1 3.2 2.9 1.8E5 15 1.4
PINN III L 500 0 98 1.6 17 0.21 0.37 51 0.41
PINN II L 50000 10 101 1.4 2.7 2.2 4.1E14 8.9 0.80
PINN II L 500 0 102 0.70 5.5 0.22 5.0E12 26 0.20
POLY L 110 6.5 16 3.5 0.35 23 2.3
MASCONS S 50000 0 119 3.1 4.4 2.8 2.7 41 8.2
PINN II S 500 0 122 2.5 8.2 2.9 3.2E5 30 3.0
MASCONS S 50000 10 128 3.3 4.6 3.0 3.0 42 9.0
MASCONS S 500 0 137 2.7 12 1.9 1.8 150 6.4
PINN I L 50000 0 140 10 4.2 8.3 1.4E5 8.8 11
PINN III L 500 10 143 6.1 27 3.0 0.36 49 6.4
PINN II S 500 10 150 5.8 12 6.6 4.3E5 32 2.5
PINN III S 500 10 157 11 58 3.4 0.31 82 22
PINN I L 50000 10 161 14 6.1 30 1.8E5 17 13
PINN I S 50000 0 165 14 8.3 38 1.3E5 31 25
PINN I S 50000 10 169 16 10 38 1.4E5 33 4.8
PINN II L 500 10 180 5.9 13 5.0 4.3E11 32 7.1
PINN I L 500 10 183 31 17 49 45000 33 32
SH S 500 10 189 25 160 5.0 0.35 1500 4.4
PINN I S 500 10 195 33 19 70 45000 37 36
SH S 500 0 198 26 160 5.0 0.35 1500 4.4
PINN I S 500 0 207 38 24 75 95000 42 47
PINN I L 500 0 210 49 38 100 1.1E5 38 55
SH S 50000 10 213 240 2000 5.0 0.35 32000 4.4
SH S 50000 0 214 230 2000 5.0 0.35 31000 4.4
PM - 500 0 215 50 43 50 50 70 120
MASCONS S 500 10 216 34 98 29 28 480 56
PM - 500 10 219 50 43 51 51 70 120
PM - 50000 10 223 53 45 53 53 70 120
PM - 50000 0 227 53 45 53 53 70 120
SH L 50000 0 240 inf 8.4E99 5.0 0.35 3.4E117 4.4
SH L 50000 10 240 inf 8.9E99 5.0 0.35 3.5E117 4.4
SH L 500 10 245 inf 2.5E97 5.0 0.35 1.1E115 4.4
SH L 500 0 251 inf 9.8E96 5.0 0.35 2.1E115 4.4
ELM L 500 0 258 76 97 300 8.0E7 99 65
ELM S 500 0 258 74 98 250 2.0E7 99 66
ELM L 500 10 260 78 97 310 8.4E7 99 65
ELM S 500 10 260 76 98 250 2.1E7 99 66
TNN S 500 10 273 770 350 1200 66000 240 640
TNN L 500 10 278 510 1100 800 56000 880 500
TNN S 50000 0 295 610 1100 1800 1.7E5 1100 580
TNN S 500 0 301 1300 1100 2600 1.5E5 990 450
TNN L 50000 10 308 630 1100 2300 2.2E5 1200 800
TNN L 50000 0 310 550 1100 1800 4.5E5 1700 640
TNN S 50000 10 310 640 1100 2400 2.2E5 1100 760
TNN L 500 0 312 1300 1300 4800 1.6E5 1100 780
ELM L 50000 10 314 9100 990 53000 2.6E9 510 47000
ELM S 50000 10 318 9200 1000 54000 1.8E8 520 1.9E6
MASCONS L 500 10 319 4.2E11 7.7E12 1.6E8 0.58 1.7E14 inf
ELM L 50000 0 324 9500 1000 56000 2.6E9 540 51000
ELM S 50000 0 328 9700 1100 57000 1.9E8 550 2.0E6
Table 2: Gravity models ranked by evaluation metrics
Results
After the models are fit on the datasets, their performance is evaluated using the accuracy metrics from Section 4 and their values are reported in Table 2. Along every metric (column), each model is colored by rank with the best model colored purple and the worst model colored red. Cells colored black correspond to values that diverged and are automatically assigned last place. These individual ranks are then summed to produce a final model score which is used to sort the table and quantify relative performance.
Table 2 should be read as follows: The first column corresponds to the specific model that was regressed, and the subsequent two columns define the data conditions used for the regression. The $N$ column corresponds to the number of training data and the error column corresponds to the magnitude of the random of the acceleration vector (i.e. either 0% or 10%). When these cells are colored light red, that means the condition is the less desirable configuration and is used to stress model performance, whereas white corresponds to the favorable data condition.
The fourth column is the aforementioned model score which sums the metric ranks for the listed model. For example, the small PINN-III (PINN-III S) is ranked 4th on the planes metric, 5th on the interior, 5th on exterior, 3rd on extrapolation, 9th on surface, and 3rd on the trajectory metric. Added together, these sum to a model score of 29 which is the best score among the 62 tested models. The individual ranks for every model in each metric are reported separately in Appendix D Table 5 for reference.
Table 2 shows that the three highest scoring gravity models are the PINN-III S, PINN-III L, and Mascon L models respectively. Each of these models are regressed under the favorable 50,000 and 0% error data conditions, and they all maintain low error across the different metrics. The PINNs generally perform better near the surface of the asteroid, where the mascon model is known to struggle; however, the mascon model performs better in the high-altitude extrapolation regime where no training data was present.
Notably, six of the eight PINN-III configurations score in the top ten models, despite some of these models being trained under noisy or sparse data conditions. For example, the 4th best model is a PINN-III S, but this model is only trained with 500 data points. The next best model trained under these conditions is a Mascon L model which scores 8th and uses a two orders of magnitude more parameters. Similarly a PINN-III L performs the best among all models trained with 10% error on the acceleration vectors — scoring 7th across all models. Finally, a PINN-III L is also the best performer of all models trained on both sparse and noisy data, scoring in the upper 35% percentile and placing 22nd total. Table 3 sorts and ranks these models by their respective training conditions for a more compact summary of model performance and again highlights how PINN-III S and L remain the two highest scoring models across all tested conditions.
Table 2 simultaneously illustrates the accuracy and robustness of the PINN-GM-III to these stressful data conditions while also exposing the brittleness of other gravity models — including past PINN-GM generations. For example, PINN-II L has the lowest error at the surface of any gravity model, averaging 2.8% across all 200,700 facets on the shape model. However, when tested in the extrapolation regime, PINN-II diverges and ranks last of all 62 models. Similarly, the Mascon L has the highest score of all analytic models when regressed on ideal training conditions, but when trained on sparse and noisy data sets, this model does not remain robust — ranking 60th of 62.
Past machine learning models have even more difficulty, with the traditional neural networks, ELM, and PINN-I consistently scoring in the lower 50th percentile. These struggles highlight just how data intensive the original machine learning implementations can be. Even with 50,000 noiseless data points, these model are unable to converge — illustrating the need for more nuanced discussion about how and when these models can break down. While the regression strategies used to fit these models sought to remain close to the original implementations, it remains possible that different training configurations and practices could yield better performance.
N 500 50000
Error (%) 10 0 10 0
Model
PINN III L 1 3 1 2
PINN III S 3 1 2 1
PINN II S 2 5 4 4
PINN II L 4 4 5 5
MASCONS L 15 2 3 3
POLY S inf inf inf 6
MASCONS S 8 6 6 8
PINN I L 5 9 7 9
POLY L inf inf inf 7
SH S 6 7 9 11
PINN I S 7 8 8 10
PM - 9 10 10 12
SH L 10 11 11 13
ELM L 12 13 14 16
TNN S 13 14 13 14
ELM S 12 13 15 17
TNN L 14 15 12 15
Table 3: Gravity models ranked and scored via evaluation metrics
Inference Time
While Tables 2 and 3 highlight the relative accuracy of the available gravity models, another important comparison point is each model’s evaluation speed. To characterize this, 1,000 randomly distributed test points are evaluated using each gravity model. The total evaluation time is measured and an average inference time per sample is calculated and reported in Figure 7.
Figure 7: Inference time per sample with the unique number of parameters in each model overlaid on the individual bars.
For the large models, Figure 7 shows that the polyhedral model of 30,006 parameters — which corresponds to approximately 10,000 facets — is the slowest gravity model averaging over 20 ms for each evaluation. The PINN-GM-III and II are an order of magnitude faster at approximately 2 ms, followed by the mascon model at 0.8 ms. For small models, the PINN-III is the slowest at 0.9 ms, followed by the remaining machine learning gravity models. The small polyhedral model (204 parameters or 66 faces) is much faster at 0.1 ms. Finally, the spherical harmonic and point mass are the fastest, executing in less than 0.01 ms per evaluation.
These results highlight that speed is among the largest trade-offs for the PINN-GM-III, particularly in the small model regime. That said, the PINN-III models are not prohibitively slow, and their computational cost remains somewhat constant irrespective of model size. For example, PINN-III L is two orders of magnitude larger than PINN-III S, but its computational cost is only twice as large. In contrast, the large polyhedral models cost appears to scale linearly, taking approximately 100 times longer for a model that is 100 times bigger. This suggests that the PINN-III remains a strong option if robustness and accuracy are the primary goal, and users can increase the size and performance of these models with relatively little added overhead.
6 Conclusions
Scientific machine learning and physics informed neural networks offer a compelling set of tools to address the gravity field modeling problem. Rather than using prescriptive analytic gravity models prone to various limitations, PINNs can learn convenient representations of the gravitational potential while maintaining desirable physics properties and assurances. While past machine learning gravity models have offered early glimpses into the potential advantages of these numerical solutions, this greater class of model remained prone to multiple pitfalls that had yet to be exposed or addressed. This paper highlights these failure cases, and introduces design modifications within the machine learning architecture that can overcome these challenges. Taken together, these modifications form the third generation PINN gravity model, or PINN-GM-III. This model is designed to solve the problems of feature divergence, bias towards low-altitude samples, numerical instability, and extrapolation error, while also proposing a framework for fusing analytic and numerical gravity models together for enhanced modeling accuracy. While these modifications are studied exclusively on the PINN gravity model, it should be noted that many of these modifications can be applied to other machine learning solutions to enhance their own robustness and generalizability.
Beyond introducing the PINN-GM-III, this manuscript also proposes new evaluation metrics to more comprehensively assess the modeling accuracy of various analytic and numerical gravity models in different orbital regimes. When trained on data from a heterogeneous density asteroid, the PINN-GM-III is shown to achieve competitive performance over past generations and other numerical and analytic models, while also demonstrating robustness to data sparse and noisy conditions. Future work will continue to investigate design modifications that can improve model performance, with a particular emphasis on returning to large celestial bodies and investigating ways in which the high-frequency components can be learned and represented more efficiently.
7 Statements and Declarations
On behalf of all authors, the corresponding author states that there is no conflict of interest.
8 Acknowledgements
This work utilized the Alpine high performance computing resource at the University of Colorado Boulder. Alpine is jointly funded by the University of Colorado Boulder, the University of Colorado Anschutz, Colorado State University, and the National Science Foundation (award 2201538).
Appendix A PINN-GM-III Training Details
Like the PINN-GM-II, the PINN-GM-III is composed of a feed-forward multi-layer perception, preceded by a feature engineering / embedding layer. Skip connections are attached between the embedding layer and each of the hidden layers of the network, and the final layer uses linear activation functions to produce the network’s prediction of the proxy potential. Unlike the PINN-GM-II, all of the experiments tested in this paper do not make use of the multi-constraint loss function due to findings presented in Appendix B.
The default hyperparameters used to train PINN-GM-III are listed in Table 4. The network is trained using the Adam optimizer with a learning rate of $2^{-8}$. The learning rate is decayed when the validation loss plateaus for 1,500 epochs. The default batch size is set to $2^{11}$ although many of the training data sizes are less than this value, so the batch size is automatically reduced to the size of the training data set when appropriate. The networks are trained for 8,192 epochs unless otherwise specified. The network is initialized using the Xavier uniform initialization scheme glorotUnderstandingDifficultyTraining2010 , and the network activation function is GELU hendrycksGaussianErrorLinear2020 . The final layer weights are initialized to zero, which heuristically led to faster convergence and better performance.
PINN-GM-III preprocesses its training data differently than PINN-GM-II. Explicitly, the position and potential are normalized by the characteristic length $x^{\star}$ and maximum potential $U^{\star}$ respectively. Using these characteristic scalars, a time constant can be computed and used in conjunction with $x^{\star}$ to non-dimensionalize the accelerations. This manifests through:
$x=\frac{\bar{x}}{x^{\star}},\quad U=\frac{\bar{U}}{U^{\star}},\quad a=\frac{\bar{a}}{a^{\star}}$ (25)
where $x^{\star}$, $U^{\star}$, and $a^{\star}$ are the non-dimensionalization constants defined as:
$x^{\star}=R,\quad U^{\star}=\max_{i}(\bar{U}_{i}-\bar{U}_{\text{LF},i}),\quad t^{\star}=\sqrt{\frac{{x^{\star}}^{2}}{U^{\star}}},\quad a^{\star}=\frac{x^{\star}}{{t^{\star}}^{2}}$ (26)
where $R$ is the maximum radius of the celestial body, $\bar{U}_{i}$ is the true gravitational potential at the training datum at $\bm{x}_{i}$, and $U_{\text{LF}}$ is any low-fidelity potential contributions already accounted for within the PINN-GM (Section 3.5).
Hyperparameter Value
learning_rate $2^{-8}$
batch_size $2^{11}$
num_epochs $2^{13}$
optimizer Adam
activation_function GELU
Hyperparameter Value
lr_scheduler plateau
lr_patience 1,500
decay_rate 0.5
min_delta 0.001
min_lr 1e-6
Table 4: Default Set of Hyperparameters for Neural Network Training
Appendix B Ablation Studies
Neural networks can be particularly sensitive to the correct choice of hyperparameters. This section seeks to characterize the sensitivity of the PINN-GM to these core hyperparameters, as well as determine the effect of different network sizes and quantity / quality of training data. Explicitly, network depth, width, batch size, learning rate, epochs, and loss function are varied, as well as the total amount data and its quality. As before, the heterogeneous density asteroid Eros detailed in Section 4 is used to provide training data. 32,768 training data are sampled uniformly between 0-3R, and the network performance is evaluated using a mean percent error averaged across 4,096 separate validation samples for each of the proposed tests.
B.1 Network Size
The first test investigates the PINN-GM’s sensitivity to network size, varying both network depth and width. The network depth is varied between ${2,4,6,8}$ hidden layers, and the width is varied between ${8,16,32,64}$ nodes per hidden layer. The remaining hyperparameters are kept fixed at the defaults provided in Table 4. The mean percent error of the models are reported in Figure 8(a) with the corresponding model size, or total trainable parameters, overlaid on the bars.
(a) Depth vs width
(b) Batch vs Learning Rate
Figure 8: Core Hyperparameters
Figure 8(a) demonstrates that PINN-GM-III performs well across a variety of different network sizes. The smallest networks will perform worse than the largest networks given their limited modeling capacity, yet despite this, all models remain below 5% error. The smallest model of 227 trainable parameters averages at 0.9% error, whereas the largest model with 30,339 parameters achieves 0.03% error. This figure illustrates that optimal models require a minimum of 16 nodes with four hidden layers. The performance continues to improve with larger models, but only marginally. Therefore, a model of six hidden layers and 16 nodes per layer (1667 parameters) or six hidden layers and 32 nodes (5891 parameters) are recommended.
B.2 Batch Size and Learning Rate
The next experiment studies the effect of batch size and learning rate on the PINN-GM. The network size is kept fixed at six hidden layers and 32 nodes per layer, but the batch size is varied between $\{2^{7},2^{9},2^{11},2^{13}\}$ and the learning rate is varied between $\{2^{-14},2^{-12},2^{-10},2^{-8}\}$. Again, the mean percent error of 4,096 samples are evaluated and shown in Figure 8(b) with the total training time overlaid on the bars rather than parameter count.
Figure 8(b) demonstrates that there are a variety of learning rates and batch sizes that are acceptable for the PINN-GM. In general, smaller batch sizes produce more accurate models, albeit this comes with much longer training times. Regarding learning rate, higher is better. This is most likely coupled with the chosen learning rate scheduler, which decreases the learning rate by a factor of 0.5 every 1,500 epochs that the validation loss does not improve. Given this learning rate scheduler configuration, a learning rate of at least $2^{-8}$ is recommended for most applications.
B.3 Data Quantity and Epochs
Fixing the learning rate to $2^{-8}$ and the batch size to $2^{11}$, a third experiment investigates the sensitivity of the PINN-GM to quantity of training data and length of training time. A PINN-III S and PINN-III L are prepared for this test. The PINN-III S uses the two layer, eight nodes per layer network configuration (227 parameters), whereas PINN-III L uses the eight layer, 64 node configuration (30,339 parameters). Both the PINN-III S and L are trained with increasing amounts of training data ranging from $\{2^{9},2^{11},2^{13},2^{15}\}$ samples and increasing the total number of training epochs from $\{2^{9},2^{11},2^{13},2^{15}\}$. As before, the mean percent error is evaluated and presented with the length of training time in Figures 9(a) and 9(b).
(a) Small Network (S)
(b) Large Network (L)
Figure 9: Data vs Epochs
Figure 9(a) shows that the PINN-III small is capable of achieving errors as low as 1% given sufficient quantities of training data and training time. Performance does decrease as the number of samples and epochs decrease; however, all models converge to solutions with less than 5% error. Similarly, Figure 9(b) demonstrates that PINN-III L consistently benefits from additional training data and epochs, with the most accurate models reaching 0.04% error. Unlike the PINN-III S, the L models always remained below 3% error even in the low data and epoch regimes. Taken together, the results suggest that the PINN-GM-III should be trained for at minimum 8,192 epochs but can benefit from longer training if resources allow.
B.4 Data Quality and Physics Constraints
The final experiment investigates the effect of additional physics constraints on model performance. Explicitly, an added Laplacian constraint in the network loss function improved model performance for the PINN-GM-II when trained on noisy data martinSmallBodies2022 . This added constraint, however, also added considerable computational overhead to compute the second order derivative of the potential via automatic differentiation. This experiment seeks to determine if this term remains necessary given the new design modifications of the PINN-GM-III.
This experiment begins by corrupting every acceleration vector by adding 10% of their magnitude in a random direction to the truth vector as discussed in Section 5. The PINN-GM-III S and L are then trained on increasing amounts of this data ranging from $N=\{2^{9},2^{11},2^{13},2^{15}\}$. These models are trained once with the proposed loss function of Section 3.2 which only penalizes errors in the acceleration vector (PINN A)
[…]
and again with the Laplacian term added (PINN AL) or
[…]
The batch size for this experiment is increased to $2^{15}$ to complete the experiment in a reasonable amount of training time, and the performance of the PINN-III L and S models as a function of training data are provided in Figure 10.
Figure 10: Error as a function of training data and loss function
Figure 10 shows that the noise in the training data does deteriorate model performance; however even with low quantities of training data, the PINN-GM S is able to achieve errors as low as 13%, just slightly above the noise floor. As the amount of training data increases, the S model reduces to approximately 7% error. The fact that these models are capable of regressing solutions beneath the noise floor is a testament to the physics-informed nature of these models. By leveraging the known dynamics of the system, the models can better ignore parts of the training data that are fundamentally inconsistent with the physics. That said, the additional Laplacian component added to the loss function (PINN AL) does not yield a compelling advantage to warrant its added training time. Therefore, it is advised that the PINN-GM does not incorporate the additional Laplacian constraint in its loss function.
Appendix C Comments on Past Machine Learning Performance
Table 1 highlights the general performance of past and present machine learning gravity models. All values are taken from their corresponding reference, but further context is warranted as each model assessed accuracy in different ways and on different asteroids. This section aims to provide relevant details regarding these metrics for completeness.
For the Gaussian process gravity model reported in Reference 22, the number of model parameters are not explicitly reported but can be deduced. Gaussian processes are defined by their covariance matrix and kernel function. The covariance matrix scales as $\mathcal{O}(N^{2})$ and the maximum number of data points used were $N=3,600$. This suggests that the minimum number of parameters used in the model is 12,960,000. The accuracy for these models are also reported at fixed radii from the center of mass for each asteroid rather than across the full domain, so the values can be considered upper-bounds. Moreover, the model is shown to diverge at high altitudes and therefore is not valid globally.
For the extreme learning machine gravity model reported in Reference 24, the model size is determined by the fact that there are 50,000 hidden nodes in the ELM. The random weights connecting the three inputs to the 50,000 hidden nodes constitutes the first 150,000 parameters, and the weights connecting the hidden layer to the acceleration output correspond to the next 150,000 parameters. Each node in the hidden layer can also have a bias, adding another 50,000 points summing to total of 350,000 total model parameters. The asteroid modeled is 25143 Itokawa, and the error in Reference 24 is reported in terms of absolute terms rather than relative terms. Using Figure 9, it can be approximated that the relative error varies between 1% and 10%.
In Reference 30, the neural network gravity model is reported to use 512 nodes per hidden layer for 6 hidden layers. The approximate number of weights and biases can be approximated by squaring the number of nodes per hidden layer and multiplying by the total number of layers minus one, and adding the biases for each node ($\approx 512^{2}*(6-1)+512*6=1,313,792$ parameters). The paper reports 1,000,000 training data were generated and divided into an 8:2 ratio between training and testing data, yielding 800,000 training data. The asteroid investigated is also 433-Eros and the average relative error of the test set is reported as 0.35% in their Table 3.
For GeodesyNets izzoGeodesyIrregularSmall2022 , SIRENs of 9 hidden layers with 100 nodes each are used ($\approx 100^{2}*(9-1)+8*100=80,800$). Four asteroids are studied: Bennu, Churyumov-Gerasimenko, Eros, and Itokawa. In their supplementary materials (Table S4), the relative error about Eros is reported at three characteristic altitudes. At their lowest altitude, the average error is 0.571% and their highest altitude is 0.146%. In attempts to quantify error across the entire high and low altitude regime, these values are averaged for the reported value of 0.359%.
Finally for the PINN-GM-III, the average error reported for a network trained with 8 hidden layers with 16 nodes per layer. The model was trained on 4,096 data points distributed between 0-10R using the same hyperparameters specified in Appendix B, and achieved an average error on the validation set of 0.3%.
Appendix D Regression Details
In Section 5, multiple gravity models are regressed and tested on a variety of data conditions. The regression procedure for the neural networks is consistent with prior explanations; however, the remaining models require further context. This section aims to provide that context, explaining how each of the remaining gravity models are regressed on the data discussed in Section 5.
D.1 Spherical Harmonics
Spherical harmonic gravity models are regressed by solving a linear system for their $p=N(N+1)$ total Stokes coefficients, where $N$ is the degree of the model. This system takes the following form:
$\bm{a}$ $=H\bm{c}$ (29)
$H^{T}\bm{a}$ $=H^{T}H\bm{c}$ (30)
$(H^{T}H)^{-1}H^{T}\bm{a}$ $=\bm{c}$ (31)
where $\bm{c}$ are the vector of spherical harmonic coefficients $C_{lm}$ and $S_{lm}$, $\bm{a}$ are the vector of accelerations, and $H$ is the jacobian $\frac{\partial\bm{a}}{\partial\bm{c}}|_{\bm{r}}$, and $\bm{r}$ is the position vector for each test point. Direct least square solutions may be appropriate for small systems, but when many harmonics need to be regressed and/or the amount of data used to regress them increases, least squares becomes computationally infeasible. Moreover, high-degree spherical harmonic models pose numerical instabilities as the observability of the harmonics decays rapidly as $(R/r)^{l}$, often making $H^{T}H$ ill-conditioned.
Kaula’s rule — a form of ridge regression that is used to regularize the spherical harmonic coefficients — can help eliminate the ill-conditioned nature of the high-degree regression. Rather than seeking to minimize the mean squared error solution
$L(\bm{c})=\lVert\bm{a}-H\bm{c}\rVert^{2}$ (32)
ridge regression adds a penalty term to ensure the coefficients decay in magnitude for higher degree harmonics through
$L(\bm{c})=\lVert\bm{a}-H\bm{c}\rVert^{2}+\lVert\Gamma\bm{c}\rVert^{2}$ (33)
where $\Gamma$ are the regularization matrix defined through Kaula’s rule
$\Gamma_{ll}=\begin{cases}\frac{\alpha}{l^{2}}&l>0\\ 1&l=0\end{cases}$ (34)
where $\alpha$ is a user specified constant typically chosen through cross validation. The corresponding solution to the ridge regression then becomes
$(H^{T}H+\Gamma)^{-1}H^{T}\bm{a}=\bm{c}\$ (35)
which yields a solution with increasingly small spherical harmonic coefficients at high-degree. This simultaneously removes the ill-posed nature of the original regression and also provides a framework for supporting low-data regression.
While using Kaula’s rule mitigates the ill-posedness of the original regression, it remains computationally expensive to invert $(H^{T}H+\Gamma)$ for large datasets. This is solved with two strategies: 1) recursive least squares and 2) iterative coefficient regression. Recursive least square sequentially feeds in small batches of data to maintain computational tractability, while iterative coefficient regression regresses low-degree harmonics before the high-degree harmonics.
For the regression used in Section 5, recursive least squares is performed in batches of 100 position / acceleration pairs using the following recursion relationships
$K_{i+1}^{-1}$ $=K_{i}^{-1}-K_{i}^{-1}H_{k}^{T}(I+H_{k}K_{i}^{-1}H_{k}^{T})^{-1}H_{k}K_{i}^{-1}$ (36)
$\bm{c}_{k+1}$ $=\bm{c}_{k}+K_{i+1}^{-1}H_{k}^{T}(\bm{a}_{k}-H_{k}\bm{c}_{k})$ (37)
where $K_{0}^{-1}=(H_{0}^{T}H_{0}+\Gamma)^{-1}$, and the iterative coefficient regression is performed by only regressing 5,000 coefficients at a time. In addition, samples beneath the Brillouin surface $r<R$ are purposefully omitted from the regression due to the $(R/r)^{l}$ scaling in the harmonic model. For very high degree models, this term diverges when evaluated on sub-Brillouin samples, which breaks the regression.
D.2 Mascons
Like spherical harmonics, mascons regression is also a form of least squares regression. Specifically, a set of $N$ mascons are uniformly distributed through the volume of the shape model, and their corresponding masses are regressed. Ideally, the mascon regression should also include a non-linear constraint applied to all mass elements that ensures they always have a value greater than zero to ensure compliance with physics — i.e. there are no such things as negative masses. This physical compliance, however, often comes at a major loss of modeling accuracy. Therefore, Section 5 allows for negative masses to be regressed to give the model the strongest cases against the PINN-GM.
In addition to allowing negative masses, the mascon regression also requires iterative fitting when the model sizes are large and there exists much data. To do this, the mascons are distributed and fit in batches of 500 using all available data. After each batch is fit, their contribution to the acceleration vectors are removed, the next batch is randomly distributed throughout the volume, and the model fits to the acceleration residuals. This process is repeated until the total mass elements are reached.
The total number of parameters for a mascon model correspond with four parameters per mascon, corresponding to the three-component position of the mass, $\bm{r}_{k}$, and it’s associated gravitational parameter $\mu_{k}$. Therefore the total parameter count for the mascon model is $p=4N$.
D.3 Extreme Learning Machines
Extreme learning machines regression also closely resembles that of the spherical harmonic regression in that it also uses ridge regression and recursive least squares optimization. Explicitly, the ELM models are regressed by applying a fixed ridge regression matrix $\Gamma_{\text{ELM}}=\alpha\mathbb{I}$ and breaking the dataset into batches to be applied recursively using Eqs. 37. Notably, the ELM also relies on a random non-linear projection into a higher-dimensional space before the linear regression as described in detail in furfaroModelingIrregularSmall2020 . For the comparison study in this paper, all input and output data are preprocessed using a min-max transformation to $[0,1]$ and applying a sigmoid activation function at the hidden nodes. The total parameter count for these models can be computed by summing the weights from the three inputs to the hidden layer, the biases of the hidden layer, and the weights to the three outputs, totaling to: $3N+N+3N=7N$.
D.4 Polyhedral Models
The regression of polyhedral models follows an entirely unrelated process to the other gravity models, requiring image data to construct the shape model of the body using stereophotoclinometry. Given that these shape models cannot be computed directly from position and acceleration data, Section 5 instead only includes the performance of polyhedral models that are of similar parameter counts to the other models and assumes these models are perfectly regressed. The model sizes of the polyhedral models are computed by summing the positions of each vertex in the model, and the indices (stored as long integers) which identify the vertices that comprise each face. Therefore the total model size can be computed via $p=3V+\frac{3F}{2}$ where $F$ corresponds to the number of facets and $V$ corresponds to the number of vertices. The $23F/$ 2 captures the fact that long integers are half the size of 32-bit floats.
D.5 Neural Network Models
The regression of the neural network models is detailed at the beginning of Section 3. The parameter count for these models corresponds with the weights connecting the $l-1$ hidden layers, the weights connecting the inputs to the first hidden layer, the weights connecting the last hidden layer to the outputs, and the biases for all nodes. Therefore, the total parameters can be estimated with $3N+N^{2}(l-1)+N+lN$ respectively where $N$ is the width of the network and $l$ is the number of hidden layers. Note that the traditional neural networks will have three outputs, whereas PINNs only have one.
D.6 Comparison Study Model Ranks
Table 5 reported the exact value for each metric for every regressed model and their rank was expressed via color in the table. Here the exact rank is reported for greater transparency.
N Error (%) Score (Rank) Planes (Rank) Interior (Rank) Exterior (Rank) Extrap. (Rank) Surface (Rank) Traj. (Rank)
Model
PINN III S 50000 0 1 4.0 5.0 5.0 3.0 9.0 3.0
PINN III L 50000 0 2 1.0 2.0 3.0 22 2.0 1.0
MASCONS L 50000 0 3 2.0 3.0 2.0 2.0 18 14
PINN III S 500 0 4 10 17 9.0 5.0 15 8.0
PINN II S 50000 0 5 5.0 4.0 10 40 6.0 5.0
PINN II L 50000 0 6 3.0 1.0 4.0 62 1.0 2.0
PINN III L 50000 10 7 17 8.0 13 6.0 5.0 27
MASCONS L 500 0 8 7.0 20 1.0 1.0 40 15
PINN III S 50000 10 9 18 16 12 9.0 21 16
MASCONS L 50000 10 11 8.0 18 6.0 4.0 46 12
POLY S 11 22 24 20 8.0 10 10
PINN II S 50000 10 12 12 7.0 16 44 7.0 9.0
PINN III L 500 0 13 11 26 7.0 20 28 6.0
PINN II L 50000 10 14 9.0 6.0 14 61 4.0 7.0
PINN II L 500 0 15 6.0 12 8.0 60 12 4.0
POLY L 16 23 25 22 18 11 11
MASCONS S 50000 0 17 15 10 15 24 24 31
PINN II S 500 0 18 13 14 17 48 13 17
MASCONS S 50000 10 19 16 11 19 25 25 32
MASCONS S 500 0 20 14 22 11 23 38 29
PINN I L 50000 0 21 24 9.0 33 38 3.0 33
PINN III L 500 10 22 21 30 18 19 27 28
PINN II S 500 10 23 19 21 32 49 16 13
PINN III S 500 10 24 25 36 21 7.0 33 35
PINN I L 50000 10 25 26 13 35 45 8.0 34
PINN I S 50000 0 26 27 15 36 37 14 36
PINN I S 50000 10 27 28 19 37 39 20 26
PINN II L 500 10 28 20 23 31 59 17 30
PINN I L 500 10 29 31 27 38 31 19 37
SH S 500 10 30 29 42 28 14 53 23
PINN I S 500 10 31 32 28 43 32 22 38
SH S 500 0 32 30 43 30 16 54 25
PINN I S 500 0 33 34 29 44 35 26 39
PINN I L 500 0 34 35 31 45 36 23 40
SH S 50000 10 35 45 56 24 10 57 21
SH S 50000 0 36 44 57 26 12 56 19
PM - 500 0 37 36 32 39 27 32 49
MASCONS S 500 10 38 33 41 34 26 41 41
PM - 500 10 39 37 33 40 28 32 49
PM - 50000 10 40 38 34 41 29 32 49
PM - 50000 0 41 39 35 42 30 32 49
SH L 50000 0 43 inf 61 25 13 61 18
SH L 50000 10 43 inf 62 23 11 62 20
SH L 500 10 44 inf 60 27 15 59 22
SH L 500 0 45 inf 59 29 17 60 24
ELM L 500 0 47 42 38 48 53 35 42
ELM S 500 0 47 40 40 46 51 37 44
ELM L 500 10 49 43 37 49 54 34 43
ELM S 500 10 49 41 39 47 52 36 45
TNN S 500 10 50 51 44 51 34 39 54
TNN L 500 10 51 46 51 50 33 47 51
TNN S 50000 0 52 48 49 53 43 50 52
TNN S 500 0 53 52 54 56 41 48 50
TNN L 50000 10 54 49 50 54 46 52 57
TNN L 50000 0 56 47 53 52 50 55 53
TNN S 50000 10 56 50 52 55 47 51 55
TNN L 500 0 57 53 55 57 42 49 56
ELM L 50000 10 58 54 45 58 57 42 58
ELM S 50000 10 59 55 46 59 55 43 60
MASCONS L 500 10 60 58 58 62 21 58 inf
ELM L 50000 0 61 56 47 60 58 44 59
ELM S 50000 0 62 57 48 61 56 45 61
Table 5: Gravity models with individual ranks per metric reported

### v3

@@
-
+The Physics-Informed Neural Network Gravity Model Generation III
 Abstract
@@
 Be it in trajectory optimization, spacecraft rendezvous, orbit determination, or other problems, gravity often plays a significant — if not dominant — role in the system dynamics.
-The mere ubiquity of this force is a testament to its significance, yet despite this, there exist surprisingly few ways to represent this force to high accuracy.
+The ubiquity of this force is a testament to its significance, yet despite this, there exist surprisingly few ways to represent this force to high accuracy.
 The construction of high-fidelity gravity models is henceforth referred to as the gravity modeling problem, and currently there exists no universally adopted solution.
@@
 These higher frequencies introduce an $\mathcal{O}(n^{2})$ computational and memory cost martinGPGPUImplementationPines2020 , and the regression of these harmonics is especially difficult due to their rapidly fading observability.
-High-order spherical harmonic models also can diverge when evaluated within the sphere that bounds all mass elements (Brillouin sphere) due to the $(R/r)^{l}$ term in the series expansion.
+High-order spherical harmonic models also can diverge when evaluated within the sphere that bounds all mass elements (Brillouin sphere).
 While these effects are negligible for near-spherical planets or moons, they can become problematic in small-body settings where objects can exhibit highly non-spherical geometries wernerExteriorGravitationPolyhedron1997 .
@@
 ELMs furfaroModelingIrregularSmall2020 350,000 768,000 1-10%
-GeodesyNet izzoGeodesyIrregularSmall2022 80,800 500,000 0.36%
+GeodesyNet izzoGeodesyIrregularSmall2022 80,800 1,000,000 0.36%
 PINN-GM-III 2,211 4,096 0.30%
@@
 Finally, in 2022 Reference 25 introduced GeodesyNets as a candidate solution to the gravity modeling problem.
-GeodesyNets are neural density fields — close relatives to the popular neural radiance fields mildenhallNeRFRepresentingScenes2022 — which uses neural networks to learn a density function for every point in a 3D volume.
-Once learned, these density fields can be integrated to yield a gravitational acceleration or potential.
-This work is an exciting new class of machine learning model capable of achieving high accuracy; albeit the cost of the numerical integration of densities make it cumbersome for simulation or on-board purposes.
-These models also have more computationally expensive training processes than other neural network approaches.
+GeodesyNets are neural density fields — close relatives to the popular neural radiance fields mildenhallNeRFRepresentingScenes2022 — which use neural networks to learn a density function for every point in a 3D volume.
+Once learned, these density fields can be numerically integrated to yield a gravitational acceleration or a value of the potential.
+This work is an exciting new class of machine learning model capable of achieving high accuracy given sufficient quantities of data.
+These models also behave more reliably at high-altitudes as the networks are not required to predict densities beyond a unit volume; albeit the cost required to perform the numerical integration can be relatively high — requiring evaluation of 30,000 to 300,000 quadrature points per prediction izzoGeodesyIrregularSmall2022 — which can make these models cumbersome to evaluate and train.
+Finally, a key feature of these models is their ability to estimate an internal density profile of an asteroid — a capability presumed available to PINNs by evaluating Poisson’s equation inside the body — but such capabilities are not the focus of this work.
 2.3 Current Challenges
@@
 martinPINNFilter2022 .
+For this study, all data will be synthetically generated and evaluation on real data is left for future work.
 Figure 2: PINN-GM Generation III with new modifications highlighted in green
@@
 These models use the same matrix operations (i.e.
-the neural network) to predict both very large and very small values of the potential depending on the test point’s altitude.
-Unfortunately, these scales can lead to difficult numerics.
+the neural network) to predict both very large and very small values of the potential depending on the test point’s altitude, which can pose difficult numerics.
 When the altitude exceeds some critical threshold, the predictions grow too small, and the network prematurely clips their value to zero.
@@
 $\hat{U}_{\text{LF}}(r)=w_{\text{LF}}U_{\text{LF}}(r)$ (21)
-where $\hat{U}_{\text{LF}}$ is the weighted low fidelity analytic model with $w_{\text{LF}}=H(r,R^{*},k^{*})$, $R^{*}=1+e$, and $k=0.5$, where $e$ is the eccentricity of the body computed via $e=\sqrt{1-b^{2}/a^{2}}$.
+where $\hat{U}_{\text{LF}}$ is the weighted low fidelity analytic model with $w_{\text{LF}}=H(r,R^{*},k^{*})$, $R^{*}=0$, and $k=0.5$.
 This ensures that at high altitude, where the analytic approximation is most accurate, it fully contributes to the final learned solution, but at lower altitudes, the analytic contribution has a reduced weight to the final solution.
@@
 To visualize how the proposed modifications affect model accuracy, six PINN-GM are trained.
-The first model corresponds with the original PINN-GM-II, and the subsequent models sequentially add one of the modification discussed above.
+The first model corresponds with the original PINN-GM-II, and the subsequent models sequentially add the proposed modifications.
 The PINN-GM are trained with a 5,000 point dataset uniformly distributed between the surface of the asteroid Eros to an altitude of 15R.
@@
 The PINN-GM-II had errors approaching 1% at the high 15R altitudes; however this decreases to 0.2%, 0.05%, 0.01%, 0.03%, and 0.001% for each of the respective modifications.
-Notably the transition to the new loss function (II) does decrease performance near the surface, but this is to be expected given the redistribution of modeling priorities that now balance accuracy at both low- and high-altitudes.
+Notably the transition to the new loss function (II) does reduce accuracy near the surface, but this is to be expected given the redistribution of modeling priorities that now balance accuracy at both low- and high-altitudes.
 Figure 4(b) illustrates the stabilizing effect of these modifications on extrapolation error — quantifying prediction error outside of the training bounds out to 100R.
-As is shown in blue, the PINN-GM-II entirely diverges after leaving the bounds of the training data at 15R; however, the added design modifications stabilize performance and prevent divergence.
-As was expected, the error in the high-altitude limit hits a numerical plateau when only the features (I) and percent loss (II) modifications are included; however, when the proxy potential (III) is used, that plateau lowers by over an order-of-magnitude.
-The boundary condition modification (IV) eliminates this plateau — albeit inducing a slight penalty near the transition point due to changes in the gradient — however, the addition of the low-fidelity point mass potential (V) eliminates this penalty and further reduces the error within the bounds of the training data.
-It should be noted that there exists a fundamental numerical limit to the percent error at approximately 0.001% where the differences between model prediction and the ground truth exists at the limit of machine precision.
+As is shown in blue, the PINN-GM-II entirely diverges after leaving the bounds of the training data at 15R; however, the design modifications stabilize performance and prevent this divergence.
+When only the features (I) and percent loss (II) modifications are included, the error in the high-altitude limit hits a numerical plateau as expected; however, the inclusion of the proxy potential (III) lowers that plateau by over an order-of-magnitude.
+The boundary condition modification (IV) eliminates this plateau — albeit inducing a small penalty near the transition point due to changes in the gradient.
+Finally, the addition of the low-fidelity point mass potential (V) eliminates this penalty and further reduces the error at higher within the bounds of the training data.
 4 Benchmarking Suite
-While the proposed modifications yield improvements as presented in Section 3.6, it remains important to characterize these gravity models from many different perspectives.
-The lack of comprehensive and standard performance benchmarks for these models are part of the reason these failure cases had not yet been identified.
+The lack of comprehensive and standard performance benchmarks for gravity models are part of the reason these failure cases had not yet been identified.
 To eliminate this possibility moving forward and to further scrutinize the proposed PINN-GM-III, six new evaluation metrics are proposed.
@@
 The second, third, and fourth metrics investigate the generalization of the model across a range of altitudes both within and beyond the training bounds.
-Explicitly, the mean acceleration error is evaluated as a function of altitude, and divided into three testing regimes: interior, exterior, and extrapolation.
+Explicitly, the mean acceleration error is evaluated as a function of altitude and divided into three testing regimes: interior, exterior, and extrapolation.
 The interior metric assesses error within the bounding sphere of radius R.
@@
 4.2 Experiment
-These metrics are used to evaluate of a PINN-GM-III trained on data generated from a synthetic, heterogeneous density asteroid modeled after 433-Eros.
+These metrics are used to evaluate a PINN-GM-III trained on data generated from a synthetic, heterogeneous density asteroid modeled after 433-Eros.
 Heterogeneous asteroids provide an especially challenging scenario for gravity models, as their internal density distributions are not directly observable.
 Some asteroids contain over- and under-dense regions within their interior, or may have been formed by two asteroids merging together.
-While some heuristic methods have been proposed to estimate these asteroids’ internal densities takahashiMorphologyDrivenDensity2014 which can be used in heterogeneous forms of the polyhedral model takahashiMorphologyDrivenDensity2014 ; the more common practice is to simply proceed with a constant density assumption.
-To induce this heterogeneous density body, two small mass inhomogeneities are placed inside the asteroid 433-Eros.
+While some heuristic methods have been proposed to estimate these asteroids’ internal densities — which can then be used in heterogeneous forms of the polyhedral model takahashiMorphologyDrivenDensity2014 — the more common practice is to simply proceed with a constant density assumption.
+To induce this heterogeneous density body, two small mass inhomogeneities are placed inside the asteroid.
 In one hemisphere, a mass element is added, and in the other hemisphere, a mass element is removed.
@@
 An additional 200,700 points are sampled on the surface of the asteroid — corresponding to a data point on every facet of the asteroid shape model.
-Together these approximately $300,000$ data points constitute a “best case” training set for the PINN-GM-III.
-These conditions are purposefully used to evaluate the upper-bound performance of the model.
-The next section will explore model performance under more realistic and stressful data conditions.
-Using this best case data, a PINN-GM-II and a PINN-GM-III are trained.
+Together these approximately $300,000$ data points constitute a “best case” training set for the PINN-GM-III used to evaluate the upper-bound performance of the model.
+Subsequent sections explore model performance under more realistic and stressful data conditions.
+Using this data, a PINN-GM-II and a PINN-GM-III are trained.
 Both networks have six hidden layers with 32 nodes per layer which corresponds to approximately 5,300 learnable parameters.
-The default hyperparameters used to train these models can be found in Appendix A alongside the ablation studies used to select them in Appendix B.
+The default hyperparameters used to train these models can be found in Appendix A alongside the studies used to select them in Appendix B.
 Once trained, these PINN-GMs are evaluated using the aforementioned metrics and compared to a constant-density polyhedral model as shown in Figure 6.
@@
 The generalization metrics are shown at the top of Figure 6, highlighting model performance across the three altitude regimes.
-The constant density polyhedral model produces the highest error both on the interior ($<$R) and exterior (1-10R) metrics with errors averaging around 10% within Brillouin sphere and above 1% error outside of it.
-In the extrapolation regime, the polyhedral model error slowly decreases as the model tends towards a point mass approximation.
-In comparison, the PINN-GM-II produced average errors of 0.5% inside the bounding sphere, and 1% outside the bounding sphere, with errors slowly increasing at higher altitudes as a result of its original absolute error loss function.
-In the extrapolation regime, the PINN-GM-II entirely diverges for the reasons discussed in Section 3.
-In contrast to these models, the PINN-GM-III maintains lower errors across the interior and exterior distributions, averaging less than 0.5% error within the bounding sphere and approximately 0.005% error outside of it.
-As the model transitions into the extrapolation regime, the Heaviside function kicks in to enforce the point mass boundary condition — helping the network converge to the same limit as the polyhedral model.
+The constant density polyhedral model produces the highest error on the interior and exterior metrics, averaging 10% and 1% respectively.
+In the extrapolation metric, the polyhedral error decreases as the model begins to behave like a point mass approximation.
+In comparison, the PINN-GM-II produces lower average errors on the interior and exterior metrics — 0.5% and 1% — however the candidate pitfalls described in Section 3 also become apparent.
+The former RMS loss function yields a monotonically increasing error as a function of altitude within the bounds of the training data, and the model diverges outside of it.
+In contrast, the PINN-GM-III maintains the lowest errors across all three regimes, averaging less than 0.5% and 0.005% error on the interior and exterior metrics, and maintains stability in the extrapolation regime due to the hyperbolic tangent function which enforces the point mass boundary conditions.
 The planes and surface metrics are shown in the middle rows of Figure 6.
-The polyhedral gravity model produced 6.5% average acceleration error on the planes, the PINN-GM-II reduces this to 0.4%, and the PINN-GM-III to 0.07%.
-The errors are largest near the surface for models, averaging at 23% for the constant density polyhedral model, 0.2% for the PINN-GM-II, and 0.18% for the PINN-GM-III.
+The polyhedral gravity model produced an average acceleration error of 6.5% on the planes, the PINN-GM-II reduces this to 0.4%, and the PINN-GM-III to 0.07%.
+The errors are largest near the surface for all models, averaging at 23% for the constant density polyhedral model, 0.2% for the PINN-GM-II, and 0.18% for the PINN-GM-III.
 Finally, the trajectory experiment is shown in the bottom row of Figure 6.
@@
 5 Comparative Study
-In this section, a comparative study is performed to evaluate the PINN-GM-III against other popular gravity models.
-Explicitly, a point mass (PM), spherical harmonic (SH), mascon, polyhedral, extreme learning machine (ELM), traditional neural network (TNN), PINN-GM-I, and PINN-GM-II model are each regressed and then evaluated using the aforementioned metrics.
+A comparative study is performed to evaluate the PINN-GM-III against other popular gravity models.
+Explicitly, a point mass (PM), spherical harmonic (SH), mascon, polyhedral, extreme learning machine (ELM), traditional neural network (TNN), PINN-GM-I, PINN-GM-II, and geodesyNet model are each regressed and then evaluated using the aforementioned metrics.
 Each model is fit eight times, permuting the data quantity, data uncertainty, and the size of each model.
@@
 Model
-PINN III S 50000 0 29 0.36 2.6 0.060 0.13 18 0.056
-PINN III L 50000 0 31 1.2E-2 0.11 1.2E-3 1.1 3.3 8.4E-3
-MASCONS L 50000 0 41 1.2E-2 0.13 5.4E-4 3.4E-4 33 2.7
-PINN III S 500 0 64 1.5 8.6 0.35 0.27 31 1.1
-PINN II S 50000 0 70 0.58 1.7 0.63 1.5E5 13 0.27
-PINN II L 50000 0 73 1.4E-2 0.059 1.4E-2 1.4E15 2.8 9.6E-3
-PINN III L 50000 10 76 3.4 3.7 2.0 0.28 13 6.3
-MASCONS L 500 0 84 0.74 11 2.1E-4 9.8E-8 310 2.7
-PINN III S 50000 10 92 3.9 8.3 2.0 0.34 35 3.0
-MASCONS L 50000 10 94 1.3 9.3 0.16 0.15 630 2.4
-POLY S 94 6.2 16 3.3 0.34 23 1.9
-PINN II S 50000 10 95 2.1 3.2 2.9 1.8E5 15 1.4
-PINN III L 500 0 98 1.6 17 0.21 0.37 51 0.41
-PINN II L 50000 10 101 1.4 2.7 2.2 4.1E14 8.9 0.80
-PINN II L 500 0 102 0.70 5.5 0.22 5.0E12 26 0.20
-POLY L 110 6.5 16 3.5 0.35 23 2.3
-MASCONS S 50000 0 119 3.1 4.4 2.8 2.7 41 8.2
-PINN II S 500 0 122 2.5 8.2 2.9 3.2E5 30 3.0
-MASCONS S 50000 10 128 3.3 4.6 3.0 3.0 42 9.0
-MASCONS S 500 0 137 2.7 12 1.9 1.8 150 6.4
-PINN I L 50000 0 140 10 4.2 8.3 1.4E5 8.8 11
-PINN III L 500 10 143 6.1 27 3.0 0.36 49 6.4
-PINN II S 500 10 150 5.8 12 6.6 4.3E5 32 2.5
-PINN III S 500 10 157 11 58 3.4 0.31 82 22
-PINN I L 50000 10 161 14 6.1 30 1.8E5 17 13
-PINN I S 50000 0 165 14 8.3 38 1.3E5 31 25
-PINN I S 50000 10 169 16 10 38 1.4E5 33 4.8
-PINN II L 500 10 180 5.9 13 5.0 4.3E11 32 7.1
-PINN I L 500 10 183 31 17 49 45000 33 32
-SH S 500 10 189 25 160 5.0 0.35 1500 4.4
-PINN I S 500 10 195 33 19 70 45000 37 36
-SH S 500 0 198 26 160 5.0 0.35 1500 4.4
-PINN I S 500 0 207 38 24 75 95000 42 47
-PINN I L 500 0 210 49 38 100 1.1E5 38 55
-SH S 50000 10 213 240 2000 5.0 0.35 32000 4.4
-SH S 50000 0 214 230 2000 5.0 0.35 31000 4.4
-PM - 500 0 215 50 43 50 50 70 120
-MASCONS S 500 10 216 34 98 29 28 480 56
-PM - 500 10 219 50 43 51 51 70 120
-PM - 50000 10 223 53 45 53 53 70 120
-PM - 50000 0 227 53 45 53 53 70 120
-SH L 50000 0 240 inf 8.4E99 5.0 0.35 3.4E117 4.4
-SH L 50000 10 240 inf 8.9E99 5.0 0.35 3.5E117 4.4
-SH L 500 10 245 inf 2.5E97 5.0 0.35 1.1E115 4.4
-SH L 500 0 251 inf 9.8E96 5.0 0.35 2.1E115 4.4
-ELM L 500 0 258 76 97 300 8.0E7 99 65
-ELM S 500 0 258 74 98 250 2.0E7 99 66
-ELM L 500 10 260 78 97 310 8.4E7 99 65
-ELM S 500 10 260 76 98 250 2.1E7 99 66
-TNN S 500 10 273 770 350 1200 66000 240 640
-TNN L 500 10 278 510 1100 800 56000 880 500
-TNN S 50000 0 295 610 1100 1800 1.7E5 1100 580
-TNN S 500 0 301 1300 1100 2600 1.5E5 990 450
-TNN L 50000 10 308 630 1100 2300 2.2E5 1200 800
-TNN L 50000 0 310 550 1100 1800 4.5E5 1700 640
-TNN S 50000 10 310 640 1100 2400 2.2E5 1100 760
-TNN L 500 0 312 1300 1300 4800 1.6E5 1100 780
-ELM L 50000 10 314 9100 990 53000 2.6E9 510 47000
-ELM S 50000 10 318 9200 1000 54000 1.8E8 520 1.9E6
-MASCONS L 500 10 319 4.2E11 7.7E12 1.6E8 0.58 1.7E14 inf
-ELM L 50000 0 324 9500 1000 56000 2.6E9 540 51000
-ELM S 50000 0 328 9700 1100 57000 1.9E8 550 2.0E6
+MASCONS L 50000 0 14 1.2E-2 0.13 5.4E-4 3.4E-4 8.3 1.7E-3
+PINN III S 50000 0 29 0.17 1.2 2.8E-2 0.065 11 4.2E-2
+PINN III L 50000 0 34 1.3E-2 0.099 1.7E-3 1.3 3.2 2.7E-3
+PINN II L 50000 0 84 4.7E-2 0.080 0.070 3.2E14 2.5 2.8E-2
+PINN III S 500 0 88 2.2 18 0.51 0.13 35 0.89
+PINN III L 50000 10 89 3.4 3.7 2.0 0.28 13 6.3
+PINN II S 50000 0 97 0.22 0.57 0.31 2.1E7 6.2 0.11
+PINN III L 500 0 98 1.6 17 0.23 0.41 52 0.67
+MASCONS L 500 0 101 0.82 14 2.5E-4 9.8E-8 350 1.3E-7
+POLY S 110 6.2 16 3.3 0.34 23 1.9
+MASCONS L 50000 10 113 1.3 9.3 0.16 0.15 610 0.37
+PINN II L 500 0 115 0.69 5.1 0.22 3.2E12 25 0.18
+MASCONS S 50000 0 117 3.1 4.4 2.8 2.7 26 12
+POLY L 119 6.5 16 3.5 0.35 23 2.3
+PINN II L 50000 10 122 1.4 2.7 2.2 4.4E14 8.8 0.68
+MASCONS S 50000 10 124 3.3 4.6 3.0 3.0 26 13
+MASCONS S 500 0 135 2.7 12 1.9 1.8 71 6.7
+PINN III L 500 10 140 6.2 23 3.1 0.36 45 6.0
+PINN II S 50000 10 140 1.5 2.9 1.8 1.1E8 9.4 6.4
+PINN II S 500 0 152 1.8 7.8 1.7 5.4E6 32 3.2
+PINN III S 500 10 159 8.2 51 3.3 0.36 76 5.0
+PINN III S 50000 10 160 8.1 51 2.3 0.32 79 17
+PINN I L 50000 0 174 10 4.3 8.4 1.6E5 8.4 12
+PINN II L 500 10 179 5.7 13 4.8 5.6E11 32 5.4
+PINN II S 500 10 182 6.4 13 6.5 8.6E6 33 2.7
+PINN I L 50000 10 183 13 5.4 16 1.8E5 14 7.0
+SH S 50000 0 202 24 370 0.45 0.22 1900 1.1
+PINN I S 50000 10 202 14 9.5 31 1.6E5 31 20
+PINN I S 50000 0 203 16 8.6 38 1.2E5 33 17
+SH S 50000 10 211 24 320 0.53 0.29 2200 1.4
+PM - 500 0 211 50 43 50 50 70 110
+PM - 500 10 215 50 43 51 51 70 110
+SH S 500 0 217 17 340 1.2 0.86 950 0.97
+PM - 50000 10 220 53 45 53 53 70 110
+PINN I L 500 10 221 29 16 48 46000 34 22
+PM - 50000 0 224 53 45 53 53 70 110
+SH S 500 10 227 27 580 1.6 0.74 1500 2.4
+PINN I S 500 10 235 33 21 69 46000 37 31
+SH L 50000 0 239 inf 1.3E99 0.45 0.22 8.8E116 1.1
+PINN I S 500 0 242 41 24 75 93000 41 42
+GEONET L 500 0 245 83 84 83 83 87 160
+SH L 50000 10 247 inf 1.8E99 0.53 0.29 1.2E117 1.4
+MASCONS S 500 10 251 34 98 29 28 460 64
+GEONET L 50000 0 252 87 87 87 87 87 160
+GEONET L 500 10 255 87 90 86 86 93 160
+GEONET L 50000 10 260 87 87 87 87 87 160
+GEONET S 500 0 273 99 99 99 99 99 180
+SH L 500 10 275 inf 2.5E97 5.0 0.35 1.0E115 4.4
+SH L 500 0 278 inf 9.8E96 5.0 0.35 2.1E115 4.4
+PINN I L 500 0 280 52 41 100 1.1E5 47 89
+GEONET S 50000 10 281 99 99 99 99 99 180
+GEONET S 500 10 288 99 99 99 99 99 180
+GEONET S 50000 0 292 99 99 99 99 99 180
+ELM L 500 0 308 76 97 300 8.0E7 99 65
+ELM L 500 10 308 78 97 310 8.4E7 99 65
+ELM S 500 0 315 74 98 250 2.0E7 99 66
+ELM S 500 10 315 76 98 250 2.1E7 99 66
+MASCONS L 500 10 369 3.6E11 5.5E12 1.7E8 0.63 1.4E14 inf
+TNN L 500 10 408 570 1200 810 56000 1100 420
+TNN S 500 0 409 1300 1100 2600 1.5E5 990 450
+TNN S 50000 0 410 630 1100 1900 1.7E5 1100 570
+TNN L 50000 0 411 550 1100 1800 3.4E5 1600 630
+TNN S 500 10 412 770 340 1200 66000 240 630
+TNN S 50000 10 413 640 1100 2300 2.2E5 1200 720
+TNN L 50000 10 414 610 1100 2300 2.2E5 1200 800
+TNN L 500 0 415 8900 1100 6300 3.2E5 810 930
+ELM L 50000 10 416 9200 1000 54000 2.6E9 520 44000
+ELM L 50000 0 417 9600 1100 56000 2.6E9 550 49000
+ELM S 50000 10 418 9200 1000 54000 1.8E8 520 1.9E6
+ELM S 50000 0 419 9700 1100 57000 1.9E8 550 2.0E6
 Table 2: Gravity models ranked by evaluation metrics
@@
 Along every metric (column), each model is colored by rank with the best model colored purple and the worst model colored red.
-Cells colored black correspond to values that diverged and are automatically assigned last place.
+Cells colored red or black correspond to values that diverged and are automatically assigned last place (70th).
+In the case of the planes, interior, exterior, extrapolation, and surface metrics, this divergence corresponds with values that exceed 100% error and therefore are not usable.
 These individual ranks are then summed to produce a final model score which is used to sort the table and quantify relative performance.
 Table 2 should be read as follows: The first column corresponds to the specific model that was regressed, and the subsequent two columns define the data conditions used for the regression.
-The $N$ column corresponds to the number of training data and the error column corresponds to the magnitude of the random of the acceleration vector (i.e.
+The $N$ column is the number of training data and the error column is the magnitude of the acceleration vector error (i.e.
 either 0% or 10%).
-When these cells are colored light red, that means the condition is the less desirable configuration and is used to stress model performance, whereas white corresponds to the favorable data condition.
+When these cells are colored light red, that means the condition is less desirable and is used to stress model performance, whereas white corresponds to the favorable data condition.
 The fourth column is the aforementioned model score which sums the metric ranks for the listed model.
-For example, the small PINN-III (PINN-III S) is ranked 4th on the planes metric, 5th on the interior, 5th on exterior, 3rd on extrapolation, 9th on surface, and 3rd on the trajectory metric.
-Added together, these sum to a model score of 29 which is the best score among the 62 tested models.
+For example, the small PINN-III (PINN-III S) is ranked 4th on the planes metric, 5th on the interior, 4th on exterior, 3rd on extrapolation, 8th on surface, and 5rd on the trajectory metric.
+Added together, these sum to a model score of 29 which is the second best score among the 70 tested models.
 The individual ranks for every model in each metric are reported separately in Appendix D Table 6 for reference.
-Table 2 shows that the three highest scoring gravity models are the PINN-III S, PINN-III L, and Mascon L models respectively.
+Table 2 shows that the three highest scoring gravity models are the Mascon L, PINN-III S, and PINN-III L models respectively.
 Each of these models are regressed under the favorable 50,000 and 0% error data conditions, and they all maintain low error across the different metrics.
-The PINNs generally perform better near the surface of the asteroid, where the mascon model is known to struggle; however, the mascon model performs better in the high-altitude extrapolation regime where no training data was present.
-Notably, six of the eight PINN-III configurations score in the top ten models, despite some of these models being trained under noisy or sparse data conditions.
-For example, the 4th best model is a PINN-III S, but this model is only trained with 500 data points.
-The next best model trained under these conditions is a Mascon L model which scores 8th and uses a two orders of magnitude more parameters.
-Similarly a PINN-III L performs the best among all models trained with 10% error on the acceleration vectors — scoring 7th across all models.
-Finally, a PINN-III L is also the best performer of all models trained on both sparse and noisy data, scoring in the upper 35% percentile and placing 22nd total.
-Table 3 sorts and ranks these models by their respective training conditions for a more compact summary of model performance and again highlights how PINN-III S and L remain the two highest scoring models across all tested conditions.
+The PINN-III L performs better near the surface of the asteroid, where the mascon model is known to struggle; however, the mascon model performs better in the high-altitude extrapolation regime where no training data was present.
+Notably, five of the eight PINN-IIIs score in the top ten models, despite some of these models being trained under noisy or sparse data conditions.
+For example, the 5th best model is a PINN-III S, but this model is only trained with 500 data points.
+The next best model trained under these conditions is a Mascon L model which scores 9th and uses a two orders of magnitude more parameters.
+Similarly a PINN-III L performs the best among all models trained with 10% error on the acceleration vectors — scoring 6th across all models.
+Finally, a PINN-III L is also the best performer of all models trained on both sparse and noisy data, scoring in the upper 30th percentile and placing 18th overall.
+Table 3(a) sorts and ranks these models by their respective training conditions for a more compact summary of model performance and again highlights how PINN-III S and L remain the two highest scoring models across all tested conditions on a heterogenous density Eros.
 Table 2 simultaneously illustrates the accuracy and robustness of the PINN-GM-III to these stressful data conditions while also exposing the brittleness of other gravity models — including past PINN-GM generations.
-For example, PINN-II L has the lowest error at the surface of any gravity model, averaging 2.8% across all 200,700 facets on the shape model.
-However, when tested in the extrapolation regime, PINN-II diverges and ranks last of all 62 models.
-Similarly, the Mascon L has the highest score of all analytic models when regressed on ideal training conditions, but when trained on sparse and noisy data sets, this model does not remain robust — ranking 60th of 62.
-Past machine learning models have even more difficulty, with the traditional neural networks, ELM, and PINN-I consistently scoring in the lower 50th percentile.
-These struggles highlight just how data intensive the original machine learning implementations can be.
-Even with 50,000 noiseless data points, these model are unable to converge — illustrating the need for more nuanced discussion about how and when these models can break down.
-While the regression strategies used to fit these models sought to remain close to the original implementations, it remains possible that different training configurations and practices could yield better performance.
+For example, PINN-II L has the lowest error at the surface of any gravity model, averaging 2.5% across all 200,700 facets on the shape model.
+However, when tested in the extrapolation regime, PINN-II diverges.
+Similarly, the Mascon L has the highest score across all models when regressed on ideal training conditions, but when trained on sparse and noisy data sets, this model diverges in all but one metric, ranking 58th of 70 overall.
+Past machine learning models exhibit greater concern, with the traditional neural networks, ELMs, PINN-I, and GeodesyNets consistently scoring in the lower 50th percentile.
+The candidate failure modes discussed in Section 3 are best captured by inspecting the interior, exterior, and extrapolation metrics.
+Note how these models exhibit deteriorating accuracy when evaluated on increasingly high altitude data — i.e.
+the interior values are generally lower than exterior values.
+Moreover, these vast majority of models — GeodesyNets excluded — meet the divergence criteria when tested beyond the bounds of the training data.
+These behaviors provide confirmation of the challenges many machine models face when tested across a more diverse set of experiments, and the PINN-GM-III avoids these same challenges due to the new suite of proposed modifications.
+As a caveat, GeodesyNets are not prone to the exact same failure modes as the other machine models due to their fundamentally different architecture and evaluation procedure.
+Because the densities predicted are only computed within a unit volume, these models do not pose the same risk of feature or model divergence at high altitudes.
+That said, these models still struggle to converge as there does not exist sufficient amounts of data to reliably constraint the density function.
+Even with 50,000 noiseless data points, the GeodesyNets yield errors in excess of 80% across the different testing regimes.
+These results further highlight how sensitive and data intensive these alternative ML model implementations can be — illustrating the need for more nuanced discussion about how and when these models can break down.
+While the regression strategies used to fit these models sought to remain close to the original implementations — as is discussed in Appendix D — it remains possible that different training configurations and practices could yield better performance.
+To ensure the conclusions drawn through this study are generalizable, three additional gravity fields are tested.
+These analyses regress the same 70 models on a homogeneous density version of Eros, and a homo- and a heterogenous density version of the asteroid Bennu.
+The imposed heterogenous density profiles again remain ±$\pm 10\%$ mass at locations $\pm$±R$/2$ along the x-axis of the respective asteroid.
+These four environments attempt to span common and exotic gravity fields: from nearly spherical to irregular geometries, and from homogeneous to heterogeneous densities profiles.
+By studying the performance across these different cases, the general behaviors of these models can be reasonably inferred for other candidate small bodies.
+The exact metric values for these additional cases are reported in Appendix D.7, and their score summaries across the different data conditions are reported in Table 3.
+Across these three additional experiments, the PINN-GM-III remains competitive — consistently reporting the highest score among the machine learning models.
+Notably, the new scenarios do show the strength of the analytic models, such as the spherical harmonic model for the asteroid Bennu, and the polyhedral model for the constant density cases.
+Importantly, the results for the polyhedral model are not directly comparable to the other models reported, as the polyhedral shape was not constructed / regressed from the position and acceleration data.
+Its inclusion can only be used to infer the strength of the model in a parametric sense, where the polyhedral models of approximately 30,000 parameters achieves better performance than other models of the same size.
+Regardless, these results imply that the PINN-GM-III exhibits its greatest utility when modeling asteroids with irregular shapes and unknown density.
 N 500 50000
@@
 Model
-PINN III L 1 3 1 2
-PINN III S 3 1 2 1
-PINN II S 2 5 4 4
-PINN II L 4 4 5 5
-MASCONS L 15 2 3 3
+PINN III L 1 2 1 3
+PINN III S 2 1 6 2
+PINN II L 3 4 3 4
+MASCONS L 15 3 2 1
+PINN II S 4 6 5 5
+MASCONS S 9 5 4 7
 POLY S inf inf inf 6
-MASCONS S 8 6 6 8
-PINN I L 5 9 7 9
-POLY L inf inf inf 7
-SH S 6 7 9 11
-PINN I S 7 8 8 10
-PM - 9 10 10 12
-SH L 10 11 11 13
-ELM L 12 13 14 16
-TNN S 13 14 13 14
-ELM S 12 13 15 17
-TNN L 14 15 12 15
-Table 3: Gravity models ranked and scored via evaluation metrics
+PINN I L 6 13 7 9
+PM - 5 7 10 12
+POLY L inf inf inf 8
+SH S 7 8 9 10
+PINN I S 8 9 8 11
+GEONET L 10 10 12 14
+SH L 11 12 11 13
+GEONET S 12 11 13 15
+TNN S 17 16 14 16
+ELM L 13 14 16 18
+ELM S 14 15 17 19
+TNN L 16 17 15 17
+Eros Heterogeneous
+N 500 50000
+Error (%) 10 0 10 0
+Model
+POLY L inf inf inf 1
+PINN III S 2 2 1 4
+PINN III L 2 2 2 3
+PINN II L 3 4 5 6
+POLY S inf inf inf 5
+PINN II S 4 6 6 7
+MASCONS L 15 3 3 2
+MASCONS S 8 5 4 8
+PM - 5 7 9 12
+PINN I L 7 13 7 9
+PINN I S 9 11 8 10
+SH S 7 8 10 11
+GEONET L 10 9 11 13
+GEONET S 12 11 13 15
+SH L 11 12 12 14
+TNN S 16 16 14 17
+ELM L 13 14 16 18
+ELM S 14 15 17 19
+TNN L 17 17 15 16
+Eros Homogeneous
+N 500 50000
+Error (%) 10 0 10 0
+Model
+SH S 1 1 1 4
+PINN III L 2 2 4 2
+PINN III S 5 3 7 3
+PINN II L 3 4 6 5
+PINN II S 6 6 3 6
+MASCONS L 15 5 5 1
+PM - 4 7 9 11
+MASCONS S 14 10 2 7
+PINN I L 7 8 8 10
+POLY L inf inf inf 8
+PINN I S 9 12 10 12
+POLY S inf inf inf 10
+SH L 8 9 11 13
+GEONET L 10 11 12 14
+GEONET S 11 13 13 15
+TNN L 16 17 14 16
+ELM L 12 15 16 18
+ELM S 14 14 17 19
+TNN S 17 16 15 17
+Bennu Heterogeneous
+N 500 50000
+Error (%) 10 0 10 0
+Model
+POLY L inf inf inf 1
+SH S 1 1 1 6
+PINN III S 4 4 2 4
+PINN III L 4 2 4 4
+POLY S inf inf inf 5
+PINN II L 5 5 6 7
+PINN II S 6 6 5 8
+MASCONS L 15 3 7 2
+PM - 2 7 8 11
+PINN I L 8 8 9 9
+MASCONS S 13 10 3 10
+PINN I S 9 12 10 12
+SH L 7 9 12 13
+GEONET L 10 11 11 14
+GEONET S 11 13 13 15
+ELM L 12 14 16 18
+ELM S 14 15 17 19
+TNN L 17 16 14 17
+TNN S 16 17 15 16
+Bennu Homogeneous
+Table 3: Summarized rank for all models across four test asteroids
 Inference Time
-While Tables 2 and 3 highlight the relative accuracy of the available gravity models, another important comparison point is each model’s evaluation speed.
+While Tables 2 and 3(a) highlight the relative accuracy of the available gravity models, another important comparison point is each model’s evaluation speed.
 To characterize this, 1,000 randomly distributed test points are evaluated using each gravity model.
-The total evaluation time is measured and an average inference time per sample is calculated and reported in Figure 7.
+The total evaluation time is measured, and an average inference time per sample is reported in Figure 7.
 Figure 7: Inference time per sample with the unique number of parameters in each model overlaid on the individual bars.
-For the large models, Figure 7 shows that the polyhedral model of 30,006 parameters — which corresponds to approximately 10,000 facets — is the slowest gravity model averaging over 20 ms for each evaluation.
-The PINN-GM-III and II are an order of magnitude faster at approximately 2 ms, followed by the mascon model at 0.8 ms.
-For small models, the PINN-III is the slowest at 0.9 ms, followed by the remaining machine learning gravity models.
-The small polyhedral model (204 parameters or 66 faces) is much faster at 0.1 ms.
-Finally, the spherical harmonic and point mass are the fastest, executing in less than 0.01 ms per evaluation.
+For the large models, Figure 7 shows that the large GeodesyNet (33,601 parameters) and the large polyhedral model (30,006 parameters / approximately 10,000 facets) are the slowest gravity models reporting approximately 40 and 10 ms for each evaluation respectively.
+The neural network models (PINN I-III and TNN) execute an order of magnitude faster, executing between 0.4 - 0.7 ms, followed by the mascon, spherical harmonic, and ELM models at 0.3 ms, 0.09 ms, and 0.08 ms respectively.
+For small models, the GeodesyNet is the slowest at 3 ms, followed by the remaining machine learning gravity models averaging between 0.3 and 0.6 ms.
+The small polyhedral model (204 parameters or 66 faces) is faster at 0.09 ms, and the mascon, spherical harmonic, and point mass models are the fastest, executing in less than 0.01 ms per evaluation.
 These results highlight that speed is among the largest trade-offs for the PINN-GM-III, particularly in the small model regime.
-That said, the PINN-III models are not prohibitively slow, and their computational cost remains somewhat constant irrespective of model size.
+That said, the PINN-III models are not prohibitively slow, and their computational cost remains nearly constant irrespective of model size.
 For example, PINN-III L is two orders of magnitude larger than PINN-III S, but its computational cost is only twice as large.
@@
 Rather than using prescriptive analytic gravity models prone to various limitations, PINNs can learn convenient representations of the gravitational potential while maintaining desirable physics properties and assurances.
-While past machine learning gravity models have offered early glimpses into the potential advantages of these numerical solutions, this greater class of model remained prone to multiple pitfalls that had yet to be exposed or addressed.
-This paper highlights these failure cases, and introduces design modifications within the machine learning architecture that can overcome these challenges.
+While past machine learning gravity models have offered early glimpses into the potential advantages of these numerical solutions, this greater class of model are often susceptible to various pitfalls that had yet to be exposed or addressed.
+This paper highlights these failure cases for a variety of past models, and introduces design modifications within the machine learning architecture that can overcome these challenges.
 Taken together, these modifications form the third generation PINN gravity model, or PINN-GM-III.
 This model is designed to solve the problems of feature divergence, bias towards low-altitude samples, numerical instability, and extrapolation error, while also proposing a framework for fusing analytic and numerical gravity models together for enhanced modeling accuracy.
-While these modifications are studied exclusively on the PINN gravity model, it should be noted that many of these modifications can be applied to other machine learning solutions to enhance their own robustness and generalizability.
+While these modifications are studied exclusively on the PINN gravity model, it should be noted that many of these modifications can likely be applied to other machine learning solutions to enhance their own robustness and generalizability.
 Beyond introducing the PINN-GM-III, this manuscript also proposes new evaluation metrics to more comprehensively assess the modeling accuracy of various analytic and numerical gravity models in different orbital regimes.
-When trained on data from a heterogeneous density asteroid, the PINN-GM-III is shown to achieve competitive performance over past generations and other numerical and analytic models, while also demonstrating robustness to data sparse and noisy conditions.
+When trained on data from both homogeneous and heterogeneous density asteroids, the PINN-GM-III is shown to achieve competitive performance over past generations and other numerical and analytic models, while also demonstrating robustness to data sparse and noisy conditions.
 Future work will continue to investigate design modifications that can improve model performance, with a particular emphasis on returning to large celestial bodies and investigating ways in which the high-frequency components can be learned and represented more efficiently.
@@
 Table 4: Default Set of Hyperparameters for Neural Network Training
-Appendix B Ablation Studies
+Appendix B Hyperparameter Optimization
 Neural networks can be particularly sensitive to the correct choice of hyperparameters.
@@
 In attempts to quantify error across the entire high and low altitude regime, these values are averaged for the reported value of 0.359%.
+The number of training data referenced is a result of the original paper sampling 1,000 data points every 10 iterations where the models were trained for 10,000 iterations, thereby equating to 1,000,000 training data.
 Finally for the PINN-GM-III, the average error reported for a network trained with 8 hidden layers with 16 nodes per layer.
@@
 Note that the traditional neural networks will have three outputs, whereas PINNs only have one.
-D.6 Comparison Study Model Ranks
-Table 5 reported the exact value for each metric for every regressed model and their rank was expressed via color in the table. Here the exact rank is reported for greater transparency.
+D.6 GeodesyNet
+The regression of the GeodesyNets follows the direct training method outlined in Ref.
+izzoGeodesyIrregularSmall2022 .
+The direct training method is chosen over the differential training method because the former does not assume knowledge of the asteroid shape — therefore matching the data conditions for the other machine learning gravity models.
+For their regression, the GeodesyNets use 300,000 quadrature points to numerically evaluate the acceleration at a field point.
+In the original paper, the models are trained for 10,000 iterations, where 1,000 new training data are randomly generated in the unit volume — i.e.
+near the body — once every 10 iterations.
+In this manuscript, the data is generated a priori — either 500 or 50,000 points spanning 0-10R — and 1,000 data points are randomly sampled from this generated data once every 10 iterations to mimic the original training configuration.
+The architecture of the GeodesyNet matched that of the other machine learning models.
+The small GeodesyNet had four hidden layers and eight nodes per hidden layer, and the large GeodesyNet had eight hidden layers with 64 nodes per hidden layer.
+All training data were non-dimensionalized to ensure that all of the asteroid density remains within the unit cube.
+The performance of the GeodesyNets in the comparative study is believed to be the result of the small dataset size and its distribution across the training volume.
+As discussed, the original GeodesyNet training paper achieves high accuracy by training on 1,000,000 data that exist within the unit cube or sphere; however, the data used in the comparative study experiments span a spherical volume out to 10R.
+This yields a volume that is approximately 1,000 times larger (volume scales cubically) with considerably less data occupying that volume.
+As such, the GeodesyNets do not have enough data to regress an accurate density field across the entire unit volume.
+D.7 Supplemental Comparison Study Results
+The following tables report the exact metric values used for the comparison study, as well as the associated rank across all models.
+D.7.1 Eros Heterogeneous Density
+N Error (%) Score Planes (%) Interior (%) Exterior (%) Extrap. (%) Surface (%) Traj. (km)
+Model
+MASCONS L 50000 0 14 1.2E-2 0.13 5.4E-4 3.4E-4 8.3 1.7E-3
+PINN III S 50000 0 29 0.17 1.2 2.8E-2 0.065 11 4.2E-2
+PINN III L 50000 0 34 1.3E-2 0.099 1.7E-3 1.3 3.2 2.7E-3
+PINN II L 50000 0 84 4.7E-2 0.080 0.070 3.2E14 2.5 2.8E-2
+PINN III S 500 0 88 2.2 18 0.51 0.13 35 0.89
+PINN III L 50000 10 89 3.4 3.7 2.0 0.28 13 6.3
+PINN II S 50000 0 97 0.22 0.57 0.31 2.1E7 6.2 0.11
+PINN III L 500 0 98 1.6 17 0.23 0.41 52 0.67
+MASCONS L 500 0 101 0.82 14 2.5E-4 9.8E-8 350 1.3E-7
+POLY S 110 6.2 16 3.3 0.34 23 1.9
+MASCONS L 50000 10 113 1.3 9.3 0.16 0.15 610 0.37
+PINN II L 500 0 115 0.69 5.1 0.22 3.2E12 25 0.18
+MASCONS S 50000 0 117 3.1 4.4 2.8 2.7 26 12
+POLY L 119 6.5 16 3.5 0.35 23 2.3
+PINN II L 50000 10 122 1.4 2.7 2.2 4.4E14 8.8 0.68
+MASCONS S 50000 10 124 3.3 4.6 3.0 3.0 26 13
+MASCONS S 500 0 135 2.7 12 1.9 1.8 71 6.7
+PINN III L 500 10 140 6.2 23 3.1 0.36 45 6.0
+PINN II S 50000 10 140 1.5 2.9 1.8 1.1E8 9.4 6.4
+PINN II S 500 0 152 1.8 7.8 1.7 5.4E6 32 3.2
+PINN III S 500 10 159 8.2 51 3.3 0.36 76 5.0
+PINN III S 50000 10 160 8.1 51 2.3 0.32 79 17
+PINN I L 50000 0 174 10 4.3 8.4 1.6E5 8.4 12
+PINN II L 500 10 179 5.7 13 4.8 5.6E11 32 5.4
+PINN II S 500 10 182 6.4 13 6.5 8.6E6 33 2.7
+PINN I L 50000 10 183 13 5.4 16 1.8E5 14 7.0
+SH S 50000 0 202 24 370 0.45 0.22 1900 1.1
+PINN I S 50000 10 202 14 9.5 31 1.6E5 31 20
+PINN I S 50000 0 203 16 8.6 38 1.2E5 33 17
+SH S 50000 10 211 24 320 0.53 0.29 2200 1.4
+PM - 500 0 211 50 43 50 50 70 110
+PM - 500 10 215 50 43 51 51 70 110
+SH S 500 0 217 17 340 1.2 0.86 950 0.97
+PM - 50000 10 220 53 45 53 53 70 110
+PINN I L 500 10 221 29 16 48 46000 34 22
+PM - 50000 0 224 53 45 53 53 70 110
+SH S 500 10 227 27 580 1.6 0.74 1500 2.4
+PINN I S 500 10 235 33 21 69 46000 37 31
+SH L 50000 0 239 inf 1.3E99 0.45 0.22 8.8E116 1.1
+PINN I S 500 0 242 41 24 75 93000 41 42
+GEONET L 500 0 245 83 84 83 83 87 160
+SH L 50000 10 247 inf 1.8E99 0.53 0.29 1.2E117 1.4
+MASCONS S 500 10 251 34 98 29 28 460 64
+GEONET L 50000 0 252 87 87 87 87 87 160
+GEONET L 500 10 255 87 90 86 86 93 160
+GEONET L 50000 10 260 87 87 87 87 87 160
+GEONET S 500 0 273 99 99 99 99 99 180
+SH L 500 10 275 inf 2.5E97 5.0 0.35 1.0E115 4.4
+SH L 500 0 278 inf 9.8E96 5.0 0.35 2.1E115 4.4
+PINN I L 500 0 280 52 41 100 1.1E5 47 89
+GEONET S 50000 10 281 99 99 99 99 99 180
+GEONET S 500 10 288 99 99 99 99 99 180
+GEONET S 50000 0 292 99 99 99 99 99 180
+ELM L 500 0 308 76 97 300 8.0E7 99 65
+ELM L 500 10 308 78 97 310 8.4E7 99 65
+ELM S 500 0 315 74 98 250 2.0E7 99 66
+ELM S 500 10 315 76 98 250 2.1E7 99 66
+MASCONS L 500 10 369 3.6E11 5.5E12 1.7E8 0.63 1.4E14 inf
+TNN L 500 10 408 570 1200 810 56000 1100 420
+TNN S 500 0 409 1300 1100 2600 1.5E5 990 450
+TNN S 50000 0 410 630 1100 1900 1.7E5 1100 570
+TNN L 50000 0 411 550 1100 1800 3.4E5 1600 630
+TNN S 500 10 412 770 340 1200 66000 240 630
+TNN S 50000 10 413 640 1100 2300 2.2E5 1200 720
+TNN L 50000 10 414 610 1100 2300 2.2E5 1200 800
+TNN L 500 0 415 8900 1100 6300 3.2E5 810 930
+ELM L 50000 10 416 9200 1000 54000 2.6E9 520 44000
+ELM L 50000 0 417 9600 1100 56000 2.6E9 550 49000
+ELM S 50000 10 418 9200 1000 54000 1.8E8 520 1.9E6
+ELM S 50000 0 419 9700 1100 57000 1.9E8 550 2.0E6
+Table 5: Metric values for each gravity model trained on heterogeneous Eros gravity field data.
 N Error (%) Score (Rank) Planes (Rank) Interior (Rank) Exterior (Rank) Extrap. (Rank) Surface (Rank) Traj. (Rank)
 Model
-PINN III S 50000 0 1 4.0 5.0 5.0 3.0 9.0 3.0
-PINN III L 50000 0 2 1.0 2.0 3.0 22 2.0 1.0
-MASCONS L 50000 0 3 2.0 3.0 2.0 2.0 18 14
-PINN III S 500 0 4 10 17 9.0 5.0 15 8.0
-PINN II S 50000 0 5 5.0 4.0 10 40 6.0 5.0
-PINN II L 50000 0 6 3.0 1.0 4.0 62 1.0 2.0
-PINN III L 50000 10 7 17 8.0 13 6.0 5.0 27
-MASCONS L 500 0 8 7.0 20 1.0 1.0 40 15
-PINN III S 50000 10 9 18 16 12 9.0 21 16
-MASCONS L 50000 10 11 8.0 18 6.0 4.0 46 12
-POLY S 11 22 24 20 8.0 10 10
-PINN II S 50000 10 12 12 7.0 16 44 7.0 9.0
-PINN III L 500 0 13 11 26 7.0 20 28 6.0
-PINN II L 50000 10 14 9.0 6.0 14 61 4.0 7.0
-PINN II L 500 0 15 6.0 12 8.0 60 12 4.0
-POLY L 16 23 25 22 18 11 11
-MASCONS S 50000 0 17 15 10 15 24 24 31
-PINN II S 500 0 18 13 14 17 48 13 17
-MASCONS S 50000 10 19 16 11 19 25 25 32
-MASCONS S 500 0 20 14 22 11 23 38 29
-PINN I L 50000 0 21 24 9.0 33 38 3.0 33
-PINN III L 500 10 22 21 30 18 19 27 28
-PINN II S 500 10 23 19 21 32 49 16 13
-PINN III S 500 10 24 25 36 21 7.0 33 35
-PINN I L 50000 10 25 26 13 35 45 8.0 34
-PINN I S 50000 0 26 27 15 36 37 14 36
-PINN I S 50000 10 27 28 19 37 39 20 26
-PINN II L 500 10 28 20 23 31 59 17 30
-PINN I L 500 10 29 31 27 38 31 19 37
-SH S 500 10 30 29 42 28 14 53 23
-PINN I S 500 10 31 32 28 43 32 22 38
-SH S 500 0 32 30 43 30 16 54 25
-PINN I S 500 0 33 34 29 44 35 26 39
-PINN I L 500 0 34 35 31 45 36 23 40
-SH S 50000 10 35 45 56 24 10 57 21
-SH S 50000 0 36 44 57 26 12 56 19
-PM - 500 0 37 36 32 39 27 32 49
-MASCONS S 500 10 38 33 41 34 26 41 41
-PM - 500 10 39 37 33 40 28 32 49
-PM - 50000 10 40 38 34 41 29 32 49
-PM - 50000 0 41 39 35 42 30 32 49
-SH L 50000 0 43 inf 61 25 13 61 18
-SH L 50000 10 43 inf 62 23 11 62 20
-SH L 500 10 44 inf 60 27 15 59 22
-SH L 500 0 45 inf 59 29 17 60 24
-ELM L 500 0 47 42 38 48 53 35 42
-ELM S 500 0 47 40 40 46 51 37 44
-ELM L 500 10 49 43 37 49 54 34 43
-ELM S 500 10 49 41 39 47 52 36 45
-TNN S 500 10 50 51 44 51 34 39 54
-TNN L 500 10 51 46 51 50 33 47 51
-TNN S 50000 0 52 48 49 53 43 50 52
-TNN S 500 0 53 52 54 56 41 48 50
-TNN L 50000 10 54 49 50 54 46 52 57
-TNN L 50000 0 56 47 53 52 50 55 53
-TNN S 50000 10 56 50 52 55 47 51 55
-TNN L 500 0 57 53 55 57 42 49 56
-ELM L 50000 10 58 54 45 58 57 42 58
-ELM S 50000 10 59 55 46 59 55 43 60
-MASCONS L 500 10 60 58 58 62 21 58 inf
-ELM L 50000 0 61 56 47 60 58 44 59
-ELM S 50000 0 62 57 48 61 56 45 61
-Table 5: Gravity models with individual ranks per metric reported
+MASCONS L 50000 0 1 1.0 3.0 2.0 2.0 4.0 2.0
+PINN III S 50000 0 2 4.0 5.0 4.0 3.0 8.0 5.0
+PINN III L 50000 0 3 2.0 2.0 3.0 22 2.0 3.0
+PINN II L 50000 0 4 3.0 1.0 5.0 70 1.0 4.0
+PINN III S 500 0 5 13 26 12 4.0 22 11
+PINN III L 50000 10 6 17 8.0 20 8.0 9.0 27
+PINN II S 50000 0 7 5.0 4.0 9.0 70 3.0 6.0
+PINN III L 500 0 8 11 25 8.0 18 27 9.0
+MASCONS L 500 0 9 7.0 21 1.0 1.0 70 1.0
+POLY S 10 20 23 27 12 11 17
+MASCONS L 50000 10 11 8.0 16 6.0 5.0 70 8.0
+PINN II L 500 0 12 6.0 12 7.0 70 13 7.0
+MASCONS S 50000 0 13 15 10 23 24 14 31
+POLY L 14 22 24 28 15 12 18
+PINN II L 50000 10 15 9.0 6.0 21 70 6.0 10
+MASCONS S 50000 10 16 16 11 24 25 15 33
+MASCONS S 500 0 17 14 18 19 23 32 29
+PINN III L 500 10 19 19 28 25 17 25 26
+PINN II S 50000 10 19 10 7.0 18 70 7.0 28
+PINN II S 500 0 20 12 14 17 70 18 21
+PINN III S 500 10 21 24 36 26 16 33 24
+PINN III S 50000 10 22 23 35 22 11 34 35
+PINN I L 50000 0 23 25 9.0 33 70 5.0 32
+PINN II L 500 10 24 18 20 29 70 17 25
+PINN II S 500 10 25 21 19 32 70 20 20
+PINN I L 50000 10 26 26 13 34 70 10 30
+SH S 50000 0 28 30 70 11 7.0 70 14
+PINN I S 50000 10 28 27 17 36 70 16 36
+PINN I S 50000 0 29 28 15 37 70 19 34
+SH S 50000 10 31 31 70 14 10 70 16
+PM - 500 0 31 37 31 39 27 31 46
+PM - 500 10 32 38 32 40 28 30 47
+SH S 500 0 33 29 70 15 21 70 12
+PM - 50000 10 34 40 33 41 29 29 48
+PINN I L 500 10 35 33 22 38 70 21 37
+PM - 50000 0 36 41 34 42 30 28 49
+SH S 500 10 37 32 70 16 20 70 19
+PINN I S 500 10 38 34 27 43 70 23 38
+SH L 50000 0 39 70 70 10 6.0 70 13
+PINN I S 500 0 40 36 29 44 70 24 39
+GEONET L 500 0 41 46 37 45 31 36 50
+SH L 50000 10 42 70 70 13 9.0 70 15
+MASCONS S 500 10 43 35 45 35 26 70 40
+GEONET L 50000 0 44 47 38 47 33 35 52
+GEONET L 500 10 45 48 40 46 32 38 51
+GEONET L 50000 10 46 49 39 48 34 37 53
+GEONET S 500 0 47 50 46 49 35 39 54
+SH L 500 10 48 70 70 30 13 70 22
+SH L 500 0 49 70 70 31 14 70 23
+PINN I L 500 0 50 39 30 70 70 26 45
+GEONET S 50000 10 51 51 47 50 36 42 55
+GEONET S 500 10 52 52 48 51 37 44 56
+GEONET S 50000 0 53 53 49 52 38 43 57
+ELM L 500 0 55 44 42 70 70 41 41
+ELM L 500 10 55 45 41 70 70 40 42
+ELM S 500 0 57 42 44 70 70 46 43
+ELM S 500 10 57 43 43 70 70 45 44
+MASCONS L 500 10 58 70 70 70 19 70 inf
+TNN L 500 10 59 70 70 70 70 70 58
+TNN S 500 0 60 70 70 70 70 70 59
+TNN S 50000 0 61 70 70 70 70 70 60
+TNN L 50000 0 62 70 70 70 70 70 61
+TNN S 500 10 63 70 70 70 70 70 62
+TNN S 50000 10 64 70 70 70 70 70 63
+TNN L 50000 10 65 70 70 70 70 70 64
+TNN L 500 0 66 70 70 70 70 70 65
+ELM L 50000 10 67 70 70 70 70 70 66
+ELM L 50000 0 68 70 70 70 70 70 67
+ELM S 50000 10 69 70 70 70 70 70 68
+ELM S 50000 0 70 70 70 70 70 70 69
+Table 6: Rank values for each gravity model trained on heterogeneous Eros gravity field data.
+D.7.2 Eros Homogeneous Density
+N Error (%) Score Planes (%) Interior (%) Exterior (%) Extrap. (%) Surface (%) Traj. (km)
+Model
+POLY L 11 2.7E-3 1.2E-2 8.9E-4 3.1E-5 0.090 2.2E-3
+MASCONS L 50000 0 20 1.2E-2 0.11 5.2E-4 3.4E-4 8.0 1.5E-3
+PINN III L 50000 0 37 2.1E-2 0.20 3.5E-3 0.52 3.7 3.2E-2
+PINN III S 50000 0 38 0.22 1.5 3.8E-2 1.3E-2 11 0.14
+POLY S 47 0.69 2.7 0.26 1.7E-2 5.7 0.36
+PINN III S 50000 10 85 2.2 3.3 1.5 3.8E-2 11 4.1
+PINN II L 50000 0 89 4.4E-2 0.090 0.066 1.2E15 2.4 3.3E-2
+PINN III L 500 0 91 1.4 16 0.20 0.061 44 0.073
+PINN III S 500 0 91 2.0 19 0.34 1.6E-2 41 0.31
+MASCONS L 500 0 105 0.82 13 2.4E-4 9.8E-8 330 6.2E-8
+PINN III L 50000 10 105 3.3 3.5 1.9 4.1E-2 13 9.1
+PINN II S 50000 0 113 0.29 0.71 0.53 1.9E6 7.9 0.44
+MASCONS S 50000 0 125 3.0 4.3 2.7 2.7 25 11
+MASCONS L 50000 10 129 1.3 9.3 0.14 0.12 610 0.30
+MASCONS S 50000 10 132 3.2 4.4 2.9 2.8 25 11
+PINN II L 500 0 132 0.68 5.2 0.20 8.8E12 26 0.57
+PINN II L 50000 10 133 1.4 2.7 2.7 3.4E14 7.8 1.0
+MASCONS S 500 0 139 2.7 12 2.0 2.0 62 7.0
+PINN II S 50000 10 151 1.9 2.8 3.0 4.5E6 12 1.3
+PINN I L 50000 0 157 6.6 2.3 5.1 31000 6.3 2.4
+PINN II S 500 0 158 1.7 6.4 1.7 3.7E5 27 1.2
+PINN I L 50000 10 179 9.7 4.7 12 64000 9.7 6.6
+PINN III S 500 10 183 28 63 9.9 4.3E-2 89 52
+PINN III L 500 10 183 26 64 9.3 4.9E-2 89 51
+PINN II L 500 10 186 5.6 13 5.0 3.5E11 31 4.3
+PINN I S 50000 0 197 11 5.1 25 24000 23 22
+PINN I S 50000 10 197 11 5.9 18 62000 20 16
+PINN II S 500 10 199 6.3 13 6.0 3.3E6 34 11
+PM - 500 0 210 51 43 52 52 71 110
+SH S 50000 0 212 25 370 0.48 0.24 2100 1.0
+PM - 500 10 214 51 43 52 52 71 110
+PM - 50000 10 218 53 44 53 53 70 110
+PM - 50000 0 222 53 44 54 53 70 110
+SH S 500 0 222 16 330 1.3 0.94 910 1.1
+SH S 50000 10 223 25 330 0.57 0.33 2200 1.4
+PINN I L 500 10 228 38 17 69 1.1E5 37 13
+SH S 500 10 228 26 550 1.6 0.82 1500 1.3
+GEONET L 50000 10 241 58 88 55 54 96 120
+MASCONS S 500 10 243 32 88 27 27 390 59
+GEONET L 500 0 243 81 82 81 81 86 160
+GEONET L 50000 0 249 87 87 87 87 87 160
+SH L 50000 0 250 inf 1.4E99 0.48 0.24 9.8E116 1.0
+PINN I S 500 10 252 44 21 95 1.2E5 45 69
+GEONET L 500 10 260 88 92 88 88 94 170
+SH L 50000 10 262 inf 2.2E99 0.57 0.33 1.3E117 1.4
+GEONET S 500 0 268 97 97 97 97 97 180
+PINN I S 500 0 268 47 21 110 1.2E5 37 66
+SH L 500 10 270 inf 2.2E97 3.7 4.0E-2 1.1E115 6.5
+SH L 500 0 273 inf 1.3E97 3.7 4.0E-2 2.1E115 6.5
+GEONET S 50000 0 278 99 99 99 99 99 180
+GEONET S 500 10 284 99 99 99 99 99 180
+PINN I L 500 0 293 88 41 160 1.4E5 43 120
+GEONET S 50000 10 294 100 100 100 100 100 180
+ELM L 500 0 310 77 98 300 7.8E7 99 64
+ELM L 500 10 310 78 98 300 8.0E7 99 65
+ELM S 500 0 315 75 98 250 2.0E7 99 66
+ELM S 500 10 315 76 98 250 2.0E7 99 66
+MASCONS L 500 10 370 3.5E11 5.1E12 1.7E8 0.64 1.3E14 inf
+TNN S 50000 10 408 470 1100 970 90000 1000 290
+TNN L 50000 10 409 470 1100 920 81000 1100 300
+TNN S 500 0 410 900 1100 2500 1.9E5 1000 390
+TNN S 500 10 411 730 1100 2100 1.5E5 1000 490
+TNN L 50000 0 412 420 1100 450 2.7E5 1400 490
+TNN S 50000 0 413 430 1000 490 33000 1100 550
+TNN L 500 10 414 680 1000 1700 1.4E5 930 560
+TNN L 500 0 415 610 1100 1800 1.7E5 860 570
+ELM L 50000 0 416 8500 920 50000 2.5E9 460 3100
+ELM L 50000 10 417 8200 890 48000 2.5E9 450 5600
+ELM S 50000 10 418 8300 900 48000 1.7E8 450 1.6E6
+ELM S 50000 0 419 8500 930 50000 1.7E8 470 1.7E6
+Table 7: Metric values for each gravity model trained on homogeneous Eros gravity field data.
+N Error (%) Score (Rank) Planes (Rank) Interior (Rank) Exterior (Rank) Extrap. (Rank) Surface (Rank) Traj. (Rank)
+Model
+POLY L 1 1.0 1.0 3.0 2.0 1.0 3.0
+MASCONS L 50000 0 2 2.0 3.0 2.0 3.0 8.0 2.0
+PINN III L 50000 0 3 3.0 4.0 4.0 19 3.0 4.0
+PINN III S 50000 0 4 5.0 6.0 5.0 4.0 11 7.0
+POLY S 5 8.0 9.0 10 6.0 4.0 10
+PINN III S 50000 10 6 16 11 18 7.0 10 23
+PINN II L 50000 0 7 4.0 2.0 6.0 70 2.0 5.0
+PINN III L 500 0 9 12 25 9.0 13 26 6.0
+PINN III S 500 0 9 15 27 11 5.0 24 9.0
+MASCONS L 500 0 11 9.0 23 1.0 1.0 70 1.0
+PINN III L 50000 10 11 20 12 21 10 13 29
+PINN II S 50000 0 12 6.0 5.0 14 70 7.0 11
+MASCONS S 50000 0 13 18 13 24 24 16 30
+MASCONS L 50000 10 14 10 20 7.0 14 70 8.0
+MASCONS S 50000 10 16 19 14 25 25 17 32
+PINN II L 500 0 16 7.0 17 8.0 70 18 12
+PINN II L 50000 10 17 11 8.0 23 70 6.0 15
+MASCONS S 500 0 18 17 21 22 23 28 28
+PINN II S 50000 10 19 14 10 26 70 12 19
+PINN I L 50000 0 20 23 7.0 30 70 5.0 22
+PINN II S 500 0 21 13 19 20 70 19 17
+PINN I L 50000 10 22 24 15 34 70 9.0 27
+PINN III S 500 10 24 32 35 33 11 35 37
+PINN III L 500 10 24 31 36 32 12 36 36
+PINN II L 500 10 25 21 22 29 70 20 24
+PINN I S 50000 0 27 25 16 36 70 15 35
+PINN I S 50000 10 27 26 18 35 70 14 34
+PINN II S 500 10 28 22 24 31 70 21 31
+PM - 500 0 29 37 31 38 27 32 45
+SH S 50000 0 30 29 70 13 16 70 14
+PM - 500 10 31 38 32 39 28 31 46
+PM - 50000 10 32 39 33 40 29 30 47
+PM - 50000 0 34 40 34 41 30 29 48
+SH S 500 0 34 27 70 17 22 70 16
+SH S 50000 10 35 28 70 16 18 70 21
+PINN I L 500 10 37 34 26 43 70 22 33
+SH S 500 10 37 30 70 19 21 70 18
+GEONET L 50000 10 38 41 40 42 31 38 49
+MASCONS S 500 10 40 33 39 37 26 70 38
+GEONET L 500 0 40 46 37 44 32 33 51
+GEONET L 50000 0 41 47 38 45 33 34 52
+SH L 50000 0 42 70 70 12 15 70 13
+PINN I S 500 10 43 35 29 47 70 27 44
+GEONET L 500 10 44 49 41 46 34 37 53
+SH L 50000 10 45 70 70 15 17 70 20
+GEONET S 500 0 47 50 42 48 35 39 54
+PINN I S 500 0 47 36 28 70 70 23 41
+SH L 500 10 48 70 70 27 8.0 70 25
+SH L 500 0 49 70 70 28 9.0 70 26
+GEONET S 50000 0 50 51 47 49 36 40 55
+GEONET S 500 10 51 52 48 50 37 41 56
+PINN I L 500 0 52 48 30 70 70 25 50
+GEONET S 50000 10 53 53 49 51 38 46 57
+ELM L 500 0 55 44 44 70 70 43 39
+ELM L 500 10 55 45 43 70 70 42 40
+ELM S 500 0 57 42 46 70 70 45 42
+ELM S 500 10 57 43 45 70 70 44 43
+MASCONS L 500 10 58 70 70 70 20 70 inf
+TNN S 50000 10 59 70 70 70 70 70 58
+TNN L 50000 10 60 70 70 70 70 70 59
+TNN S 500 0 61 70 70 70 70 70 60
+TNN S 500 10 62 70 70 70 70 70 61
+TNN L 50000 0 63 70 70 70 70 70 62
+TNN S 50000 0 64 70 70 70 70 70 63
+TNN L 500 10 65 70 70 70 70 70 64
+TNN L 500 0 66 70 70 70 70 70 65
+ELM L 50000 0 67 70 70 70 70 70 66
+ELM L 50000 10 68 70 70 70 70 70 67
+ELM S 50000 10 69 70 70 70 70 70 68
+ELM S 50000 0 70 70 70 70 70 70 69
+Table 8: Rank values for each gravity model trained on homogeneous Eros gravity field data.
+D.7.3 Bennu Heterogeneous Density
+N Error (%) Score Planes (%) Interior (%) Exterior (%) Extrap. (%) Surface (%) Traj. (km)
+Model
+MASCONS L 50000 0 13 1.0E-2 0.37 1.4E-3 7.1E-4 3.3 5.4E-5
+PINN III L 50000 0 26 1.4E-2 0.62 1.6E-3 0.31 3.0 4.4E-5
+PINN III S 50000 0 31 0.11 1.6 3.1E-2 3.2E-2 3.3 1.9E-4
+SH S 50000 0 39 0.22 3.9 0.072 1.7E-3 6.1 1.2E-4
+SH S 50000 10 66 0.35 3.8 0.21 0.16 6.0 1.0E-2
+SH S 500 0 79 0.37 4.1 0.23 0.16 6.2 1.3E-2
+PINN II L 50000 0 93 4.5E-2 0.39 3.9E-2 3.5E16 1.4 4.0E-3
+PINN III L 500 0 94 0.49 17 0.12 0.053 28 1.8E-4
+PINN II S 50000 0 102 0.14 1.2 0.16 2.8E5 2.6 2.3E-3
+MASCONS S 50000 10 111 1.1 5.7 0.92 0.83 10 0.066
+PINN III S 500 0 113 0.72 19 0.20 0.14 32 1.9E-3
+MASCONS S 50000 0 116 1.3 5.6 1.0 0.97 10 0.075
+PINN II L 500 0 122 0.28 5.3 0.13 4.1E12 8.4 9.1E-4
+MASCONS L 500 0 123 0.42 25 1.7E-2 5.0E-8 140 2.6E-11
+PINN II S 500 0 129 0.48 4.1 0.62 2.6E7 6.2 1.8E-3
+SH S 500 10 136 2.4 15 1.3 0.81 19 0.065
+POLY L 137 6.3 14 3.5 0.35 15 9.4E-3
+PINN II S 50000 10 147 1.1 2.6 1.4 4.7E7 4.4 0.084
+POLY S 148 6.4 15 3.5 0.36 16 1.1E-2
+PINN I L 50000 0 148 1.6 1.7 1.4 1.0E5 2.8 0.12
+PINN III L 50000 10 151 2.3 14 1.6 0.33 20 0.24
+MASCONS L 50000 10 154 0.79 27 0.17 0.10 160 7.5E-3
+PINN II L 50000 10 157 1.2 4.1 1.7 3.2E13 6.1 0.072
+PINN III S 50000 10 174 4.5 19 1.9 0.33 30 0.31
+PINN I L 50000 10 180 3.8 4.0 10 83000 6.2 0.11
+PM - 50000 10 184 6.9 17 4.1 2.2 20 0.17
+PINN III L 500 10 184 5.7 22 2.7 0.35 34 0.27
+PM - 50000 0 186 6.9 17 4.2 2.3 20 0.18
+PINN II L 500 10 188 3.9 9.5 4.0 1.2E11 12 0.067
+PM - 500 0 188 6.9 17 4.2 2.4 20 0.18
+PM - 500 10 191 6.9 17 4.4 2.6 20 0.20
+PINN III S 500 10 192 7.3 20 3.0 0.34 31 0.34
+PINN I S 50000 0 194 5.2 7.7 18 83000 13 0.058
+PINN II S 500 10 200 4.5 8.7 5.4 1.3E7 11 0.14
+PINN I S 50000 10 210 6.7 8.5 17 69000 15 0.11
+PINN I L 500 0 216 5.9 7.8 11 53000 11 0.44
+SH L 50000 0 224 2300 2.0E7 0.072 1.7E-3 5.4E8 1.2E-4
+PINN I L 500 10 228 8.1 9.9 21 58000 13 0.19
+SH L 500 0 242 7.4 5100 3.7 0.37 1.8E5 9.8E-3
+MASCONS S 500 0 244 25 46 24 24 67 1.1
+SH L 500 10 244 7.8 9900 3.7 0.37 2.9E5 9.8E-3
+GEONET L 500 0 246 39 45 38 38 50 2.0
+SH L 50000 10 249 9500 3.1E7 0.21 0.16 1.9E9 1.0E-2
+GEONET L 50000 0 250 44 44 44 44 44 2.2
+PINN I S 500 10 250 12 13 43 59000 20 0.56
+PINN I S 500 0 253 13 14 47 51000 21 0.24
+GEONET L 50000 10 260 51 52 51 51 52 2.5
+GEONET L 500 10 276 85 94 84 84 95 3.3
+GEONET S 500 10 281 91 95 91 91 95 3.5
+GEONET S 500 0 289 95 96 95 95 97 3.6
+GEONET S 50000 0 301 98 98 98 98 98 3.6
+GEONET S 50000 10 309 98 98 98 98 98 3.6
+ELM S 500 0 322 62 97 220 1.8E7 98 0.21
+ELM L 500 0 324 59 97 240 4.9E7 97 0.45
+ELM L 500 10 325 59 97 240 4.9E7 97 0.44
+MASCONS S 500 10 326 87 150 84 81 210 2.7
+ELM S 500 10 326 62 97 220 1.8E7 98 0.21
+MASCONS L 500 10 358 1.9E11 1.2E13 5.8E9 0.25 5.1E13 10
+TNN S 500 0 399 400 490 840 71000 470 1.9
+TNN L 500 0 401 380 620 770 65000 610 2.2
+TNN L 50000 10 405 450 730 1200 1.8E5 670 3.3
+TNN S 50000 10 408 450 720 1200 1.1E5 730 3.5
+TNN L 50000 0 412 450 840 1300 2.3E5 870 5.5
+TNN S 50000 0 413 450 780 1300 1.2E5 770 5.5
+TNN L 500 10 414 390 540 850 73000 520 6.5
+TNN S 500 10 415 400 510 950 79000 500 6.8
+ELM L 50000 10 417 4300 600 25000 1.9E9 520 560
+ELM L 50000 0 418 4200 580 24000 1.9E9 500 670
+ELM S 50000 10 419 4500 620 26000 8.3E7 530 18000
+ELM S 50000 0 420 4300 600 25000 8.2E7 520 18000
+Table 9: Metric values for each gravity model trained on heterogeneous Bennu gravity field data.
+N Error (%) Score (Rank) Planes (Rank) Interior (Rank) Exterior (Rank) Extrap. (Rank) Surface (Rank) Traj. (Rank)
+Model
+MASCONS L 50000 0 1 1.0 1.0 1.0 2.0 5.0 3.0
+PINN III L 50000 0 2 2.0 3.0 2.0 13 4.0 2.0
+PINN III S 50000 0 3 4.0 5.0 4.0 5.0 6.0 7.0
+SH S 50000 0 4 6.0 9.0 6.0 4.0 9.0 5.0
+SH S 50000 10 5 8.0 8.0 14 10 8.0 18
+SH S 500 0 6 9.0 13 15 11 11 20
+PINN II L 50000 0 7 3.0 2.0 5.0 70 1.0 12
+PINN III L 500 0 8 12 29 8.0 6.0 33 6.0
+PINN II S 50000 0 9 5.0 4.0 10 70 2.0 11
+MASCONS S 50000 10 10 16 16 17 23 16 23
+PINN III S 500 0 11 13 34 12 8.0 36 10
+MASCONS S 50000 0 12 18 15 18 24 15 26
+PINN II L 500 0 13 7.0 14 9.0 70 14 8.0
+MASCONS L 500 0 14 10 38 3.0 1.0 70 1.0
+PINN II S 500 0 15 11 11 16 70 12 9.0
+SH S 500 10 16 21 27 19 22 25 22
+POLY L 17 29 26 27 18 23 14
+PINN II S 50000 10 18 15 7.0 21 70 7.0 27
+POLY S 20 30 28 28 19 24 19
+PINN I L 50000 0 20 19 6.0 20 70 3.0 30
+PINN III L 50000 10 21 20 25 22 14 31 39
+MASCONS L 50000 10 22 14 39 11 7.0 70 13
+PINN II L 50000 10 23 17 12 23 70 10 25
+PINN III S 50000 10 24 24 35 24 15 34 42
+PINN I L 50000 10 25 22 10 37 70 13 28
+PM - 50000 10 27 32 33 32 25 30 32
+PINN III L 500 10 27 27 37 25 17 37 41
+PM - 50000 0 28 33 32 33 26 29 33
+PINN II L 500 10 30 23 21 31 70 19 24
+PM - 500 0 30 34 31 34 27 28 34
+PM - 500 10 31 35 30 35 28 27 36
+PINN III S 500 10 32 36 36 26 16 35 43
+PINN I S 50000 0 33 26 17 40 70 20 21
+PINN II S 500 10 34 25 20 36 70 18 31
+PINN I S 50000 10 35 31 19 39 70 22 29
+PINN I L 500 0 36 28 18 38 70 17 45
+SH L 50000 0 37 70 70 7.0 3.0 70 4.0
+PINN I L 500 10 38 39 22 41 70 21 35
+SH L 500 0 39 37 70 29 20 70 16
+MASCONS S 500 0 41 42 42 42 29 41 48
+SH L 500 10 41 38 70 30 21 70 15
+GEONET L 500 0 42 43 41 43 30 39 50
+SH L 50000 10 43 70 70 13 9.0 70 17
+GEONET L 50000 0 45 44 40 45 31 38 52
+PINN I S 500 10 45 40 23 44 70 26 47
+PINN I S 500 0 46 41 24 46 70 32 40
+GEONET L 50000 10 47 45 43 47 32 40 53
+GEONET L 500 10 48 50 44 49 34 43 56
+GEONET S 500 10 49 52 45 50 35 42 57
+GEONET S 500 0 50 53 46 51 36 44 59
+GEONET S 50000 0 51 54 51 52 37 47 60
+GEONET S 50000 10 52 55 52 53 38 50 61
+ELM S 500 0 53 48 49 70 70 48 37
+ELM L 500 0 54 46 47 70 70 45 46
+ELM L 500 10 55 47 48 70 70 46 44
+MASCONS S 500 10 57 51 70 48 33 70 54
+ELM S 500 10 57 49 50 70 70 49 38
+MASCONS L 500 10 58 70 70 70 12 70 66
+TNN S 500 0 59 70 70 70 70 70 49
+TNN L 500 0 60 70 70 70 70 70 51
+TNN L 50000 10 61 70 70 70 70 70 55
+TNN S 50000 10 62 70 70 70 70 70 58
+TNN L 50000 0 63 70 70 70 70 70 62
+TNN S 50000 0 64 70 70 70 70 70 63
+TNN L 500 10 65 70 70 70 70 70 64
+TNN S 500 10 66 70 70 70 70 70 65
+ELM L 50000 10 67 70 70 70 70 70 67
+ELM L 50000 0 68 70 70 70 70 70 68
+ELM S 50000 10 69 70 70 70 70 70 69
+ELM S 50000 0 70 70 70 70 70 70 70
+Table 10: Rank values for each gravity model trained on heterogeneous Bennu gravity field data.
+D.7.4 Bennu Homogeneous Density
+N Error (%) Score Planes (%) Interior (%) Exterior (%) Extrap. (%) Surface (%) Traj. (km)
+Model
+POLY L 8 2.9E-3 3.6E-2 1.0E-3 6.9E-5 0.21 1.1E-5
+MASCONS L 50000 0 19 9.8E-3 0.34 1.5E-3 7.1E-4 3.2 5.3E-5
+PINN III S 50000 0 35 0.11 1.4 3.1E-2 1.1E-2 2.8 3.7E-4
+PINN III L 50000 0 35 1.3E-2 0.52 1.6E-3 2.2E-2 3.1 6.7E-5
+POLY S 56 0.31 2.7 0.13 9.2E-3 3.6 1.2E-3
+SH S 50000 0 57 0.18 3.4 0.061 1.1E-2 5.5 7.3E-4
+SH S 500 0 83 0.30 3.4 0.18 0.12 5.4 9.0E-3
+SH S 50000 10 89 0.34 3.4 0.21 0.16 5.4 1.1E-2
+PINN II L 50000 0 99 0.050 0.39 0.050 1.6E17 2.6 8.8E-4
+PINN II S 50000 0 101 0.11 1.0 0.11 95000 2.3 3.5E-4
+PINN III L 500 0 104 0.42 17 0.073 1.6E-2 29 3.2E-4
+MASCONS L 500 0 126 0.43 25 1.8E-2 5.0E-8 140 2.6E-11
+PINN I L 50000 0 127 0.86 1.0 0.74 20000 2.0 5.2E-3
+PINN III S 50000 10 129 1.5 3.9 1.0 2.6E-2 6.7 0.19
+PINN III S 500 0 130 1.0 41 0.18 1.0E-2 50 8.7E-4
+PINN II L 500 0 131 0.23 3.8 0.12 8.7E11 6.2 8.3E-4
+MASCONS S 50000 10 136 1.4 5.6 1.1 1.1 10 0.084
+MASCONS S 50000 0 140 1.5 5.5 1.3 1.2 10 0.094
+PINN III L 50000 10 141 2.2 6.5 1.1 1.9E-2 11 0.21
+PINN II S 500 0 142 0.50 3.0 0.60 7.2E6 5.1 1.6E-2
+PINN II S 50000 10 149 0.78 2.3 1.2 7.5E6 4.0 4.5E-2
+SH S 500 10 160 2.3 15 1.3 0.84 19 0.066
+PINN II L 50000 10 168 1.2 3.5 1.6 2.1E14 4.9 0.074
+MASCONS L 50000 10 169 0.79 27 0.17 0.11 160 7.7E-3
+PM - 500 0 169 2.6 9.8 1.8 1.4 13 0.12
+PM - 500 10 172 2.7 9.7 2.0 1.7 13 0.14
+PM - 50000 10 174 2.7 9.7 2.1 1.7 13 0.14
+PM - 50000 0 176 2.8 9.7 2.1 1.8 13 0.15
+PINN I L 50000 10 195 3.0 3.4 5.1 6500 5.5 0.10
+PINN I L 500 0 205 3.7 6.4 9.6 67000 9.3 0.070
+PINN III S 500 10 209 26 18 9.4 2.2E-2 29 0.80
+PINN III L 500 10 209 25 18 9.2 2.2E-2 29 0.79
+PINN II L 500 10 211 4.0 9.6 3.9 2.0E9 12 0.067
+PINN II S 500 10 212 3.9 9.5 4.6 7.2E6 12 0.083
+SH L 500 0 217 2.5 6800 0.82 2.0E-2 1.2E5 1.7E-2
+PINN I S 50000 10 220 4.4 5.2 8.9 5600 9.5 0.22
+SH L 500 10 223 2.9 11000 0.82 2.0E-2 2.2E5 1.7E-2
+PINN I S 50000 0 226 4.4 6.4 17 12000 11 0.21
+SH L 50000 0 234 2500 1.4E7 0.061 1.1E-2 4.7E8 7.3E-4
+MASCONS S 500 0 236 20 33 20 19 47 1.0
+PINN I L 500 10 241 9.6 10 27 81000 15 0.13
+GEONET L 500 0 244 32 40 32 32 46 1.8
+GEONET L 50000 0 246 38 38 38 38 38 2.0
+GEONET L 50000 10 259 59 59 59 59 59 2.7
+PINN I S 500 0 260 11 12 46 62000 18 0.77
+SH L 50000 10 262 8700 4.1E7 0.21 0.16 1.9E9 1.1E-2
+PINN I S 500 10 264 15 18 48 74000 25 0.37
+GEONET L 500 10 273 83 94 83 83 95 3.3
+GEONET S 500 0 285 97 98 97 97 98 3.6
+GEONET S 500 10 293 97 98 97 97 98 3.6
+GEONET S 50000 10 299 98 98 98 98 98 3.6
+GEONET S 50000 0 305 99 99 99 99 99 3.6
+ELM L 500 0 318 59 97 240 4.9E7 98 0.43
+ELM S 500 0 319 62 98 220 1.8E7 98 0.21
+ELM L 500 10 320 60 97 240 5.0E7 98 0.42
+MASCONS S 500 10 323 82 130 80 78 180 3.0
+ELM S 500 10 324 63 98 220 1.8E7 98 0.21
+MASCONS L 500 10 363 1.9E11 1.1E13 6.1E9 0.25 5.1E13 10
+TNN L 50000 10 408 450 720 360 8800 660 4.5
+TNN S 50000 10 409 380 660 280 16000 650 4.7
+TNN S 50000 0 410 380 690 340 34000 680 4.8
+TNN L 50000 0 411 380 810 320 1.0E5 820 5.2
+TNN L 500 0 413 700 560 3000 1.1E5 510 14
+TNN S 500 0 414 430 490 1000 90000 490 15
+TNN S 500 10 415 480 510 1200 1.1E5 520 15
+TNN L 500 10 416 440 550 1100 1.0E5 550 17
+ELM L 50000 0 417 4000 560 23000 1.9E9 480 36
+ELM L 50000 10 418 4200 580 24000 1.9E9 490 120
+ELM S 50000 0 419 4200 580 24000 8.3E7 490 18000
+ELM S 50000 10 420 4300 600 25000 8.4E7 510 19000
+Table 11: Metric values for each gravity model trained on homogeneous Bennu gravity field data.
+N Error (%) Score (Rank) Planes (Rank) Interior (Rank) Exterior (Rank) Extrap. (Rank) Surface (Rank) Traj. (Rank)
+Model
+POLY L 1 1.0 1.0 1.0 2.0 1.0 2.0
+MASCONS L 50000 0 2 2.0 2.0 2.0 3.0 7.0 3.0
+PINN III S 50000 0 4 5.0 7.0 5.0 6.0 5.0 7.0
+PINN III L 50000 0 4 3.0 4.0 3.0 15 6.0 4.0
+POLY S 5 10 9.0 12 4.0 8.0 13
+SH S 50000 0 6 7.0 13 7.0 7.0 14 9.0
+SH S 500 0 7 9.0 12 15 18 13 16
+SH S 50000 10 8 11 11 17 20 12 18
+PINN II L 50000 0 9 4.0 3.0 6.0 70 4.0 12
+PINN II S 50000 0 10 6.0 6.0 10 70 3.0 6.0
+PINN III L 500 0 11 12 33 9.0 9.0 36 5.0
+MASCONS L 500 0 12 13 37 4.0 1.0 70 1.0
+PINN I L 50000 0 13 17 5.0 19 70 2.0 14
+PINN III S 50000 10 14 21 17 22 16 17 36
+PINN III S 500 0 15 18 42 14 5.0 40 11
+PINN II L 500 0 16 8.0 16 11 70 16 10
+MASCONS S 50000 10 17 20 20 24 23 21 28
+MASCONS S 50000 0 18 22 19 26 24 20 29
+PINN III L 50000 10 19 23 23 23 10 23 39
+PINN II S 500 0 20 14 10 18 70 11 19
+PINN II S 50000 10 21 15 8.0 25 70 9.0 22
+SH S 500 10 22 24 32 27 22 32 23
+PINN II L 50000 10 23 19 15 28 70 10 26
+MASCONS L 50000 10 25 16 38 13 17 70 15
+PM - 500 0 25 26 29 29 25 29 31
+PM - 500 10 26 27 28 30 26 28 33
+PM - 50000 10 27 28 27 31 27 27 34
+PM - 50000 0 28 29 26 32 28 26 35
+PINN I L 50000 10 29 31 14 35 70 15 30
+PINN I L 500 0 30 32 21 39 70 18 25
+PINN III S 500 10 32 42 35 38 13 34 47
+PINN III L 500 10 32 41 36 37 14 35 46
+PINN II L 500 10 33 34 25 33 70 25 24
+PINN II S 500 10 34 33 24 34 70 24 27
+SH L 500 0 35 25 70 20 11 70 21
+PINN I S 50000 10 36 36 18 36 70 19 41
+SH L 500 10 37 30 70 21 12 70 20
+PINN I S 50000 0 38 35 22 40 70 22 37
+SH L 50000 0 39 70 70 8.0 8.0 70 8.0
+MASCONS S 500 0 40 40 39 41 29 39 48
+PINN I L 500 10 41 37 30 42 70 30 32
+GEONET L 500 0 42 43 41 43 30 38 49
+GEONET L 50000 0 43 44 40 44 31 37 50
+GEONET L 50000 10 44 45 43 47 32 41 51
+PINN I S 500 0 45 38 31 45 70 31 45
+SH L 50000 10 46 70 70 16 19 70 17
+PINN I S 500 10 47 39 34 46 70 33 42
+GEONET L 500 10 48 51 44 49 34 42 53
+GEONET S 500 0 49 52 49 50 35 45 54
+GEONET S 500 10 50 53 50 51 36 48 55
+GEONET S 50000 10 51 54 51 52 37 49 56
+GEONET S 50000 0 52 55 52 53 38 50 57
+ELM L 500 0 53 46 45 70 70 43 44
+ELM S 500 0 54 48 47 70 70 46 38
+ELM L 500 10 55 47 46 70 70 44 43
+MASCONS S 500 10 56 50 70 48 33 70 52
+ELM S 500 10 57 49 48 70 70 47 40
+MASCONS L 500 10 58 70 70 70 21 70 62
+TNN L 50000 10 59 70 70 70 70 70 58
+TNN S 50000 10 60 70 70 70 70 70 59
+TNN S 50000 0 61 70 70 70 70 70 60
+TNN L 50000 0 62 70 70 70 70 70 61
+TNN L 500 0 63 70 70 70 70 70 63
+TNN S 500 0 64 70 70 70 70 70 64
+TNN S 500 10 65 70 70 70 70 70 65
+TNN L 500 10 66 70 70 70 70 70 66
+ELM L 50000 0 67 70 70 70 70 70 67
+ELM L 50000 10 68 70 70 70 70 70 68
+ELM S 50000 0 69 70 70 70 70 70 69
+ELM S 50000 10 70 70 70 70 70 70 70
+Table 12: Rank values for each gravity model trained on homogeneous Bennu gravity field data.